### v3

# Data-driven Error Estimation: Upper Bounding Multiple Errors without Class Complexity as Input

Abstract
Constructing confidence intervals that are simultaneously valid across a class of estimates is central for tasks such as multiple mean estimation, bounding generalization error in machine learning, and adaptive experimental design. We frame this as an “error estimation problem,” where the goal is to determine a high-probability upper bound on the maximum error for a class of estimates. We propose an entirely data-driven approach that derives such bounds for both finite and infinite class settings, naturally adapting to a potentially unknown correlation structure of random errors. Notably, our method does not require class complexity as an input, overcoming a major limitation of existing approaches such as union bounding and bounds based on Talagrand’s inequality. In this paper, we present our simple yet general solution and demonstrate its flexibility through applications ranging from constructing multiple simultaneously valid confidence intervals to optimizing exploration in contextual bandit algorithms.
Machine Learning, ICML, Bandits, Generalization Bounds
1 Introduction
In statistical estimation, quantifying uncertainty involves bounding the error ($e$) between a target estimand ($\theta$) and its estimate ($\hat{\theta}$). While single errors are often bounded using concentration inequalities or the central limit theorem (CLT), many practical applications also require constructing simultaneous confidence intervals (CI) over multiple estimates. We frame this as the “error estimation problem”:
For each $h$ in a class of estimation tasks $\mathcal{H}$, we define an estimand $\theta_{h}$, and its estimate $\hat{\theta}_{h}$. The error $e_{h}$, is a measure of distance between the estimand and the estimate. Our goal is to derive a high-probability upper bound on $\max_{h\in\mathcal{H}}e_{h}$.
Consider for instance an educational assessment where $\mathcal{H}$ represents a (finite or infinite) set of various demographic or academic subgroups (such as ‘all students”,“male students”, “4th graders”, etc.). Setting $\theta_{h}$ to be the true mean test score for subgroup $h$, and $\hat{\theta}_{h}$ as its sample mean estimate, solving the error estimation problem would allow us to construct simultaneously valid CI for all subgroup mean test scores.
For finite estimation classes, a standard solution is the union bound, which adjusts individual error bounds based on the number of subgroups, $|\mathcal{H}|$. However, this method yields conservative bounds and doesn’t extend to cases with infinitely many subgroups. Another example of an infinite class arises when bounding the excess risk of a model – the gap between a learned model’s error and the error of the best model in the model class – can be framed as an error estimation problem over the model class. Existing methods for infinite classes rely on the need to characterize class complexity (e.g.VC dimension or Rademacher complexity). These measures can be hard to determine precisely, leading to the use of upper bounds which result in overly conservative CI in practice.
Motivated by these limitations, in this paper we present a new data-driven method for constructing valid confidence intervals. Our approach: (a) circumvents the need for class complexity characterization - a major limitation that, to the best of our knowledge, has not been overcome before, and (b) naturally adapts to the true correlation structure of errors, yielding much tighter bounds in cases where traditional methods are overly conservative. We achieve this goal by leveraging single error bounds, which can be either finite-sample or asymptotic, to create corresponding uniform error bounds through a surprisingly simple procedure (Theorem 2.2).
To provide intuition for our approach, we draw an analogy between our solution to bounding maximum error over a class and the well-known overfitting phenomenon that often occurs in machine learning. The goal in many learning problems is to minimize true risk (expected loss over a data distribution) over a class of predictors $\mathcal{F}$. Empirical risk minimization (ERM) typically approximates this by selecting $f\in\mathcal{F}$ to minimize empirical risk (average loss) on some dataset $S$. This inherently “overfits” to $S$: with high probability, the minimum empirical risk is less than the minimum true risk (up to a small correction). [^1] This is closely related to the fact that training error is nearly always less than test error. Formally, if we let $\mathcal{R}(f)$ be the true risk and ^$\hat{\mathcal{R}}(f)$ be the empirical risk, we see with high probability
$\min_{f\in\mathcal{F}}\hat{\mathcal{R}}(f)\lesssim\min_{f\in\mathcal{F}}\mathcal{R}(f)$ (1)
While overfitting is often viewed as an issue to be mitigated, we view it as an informative data-driven technique that can be used to construct a high probability lower bound on the minimum true risk (without ever needing to rely on class complexity bounds of $\mathcal{F}$).
Error estimation, like minimizing risk, involves optimizing an unknown quantity over a class. However, unlike fixed risks, errors are data-dependent random variables. To tackle this, we introduce a holdout dataset, that is not used for constructing our estimates and defining our errors. We then condition on the input estimates and error definitions, and are thus able to treat them as fixed quantities under the sampling distribution of the holdout dataset. Next we use the holdout dataset to construct high probability upper bounds on individual errors. Theorem 2.2, which is analogous to the overfitting inequality (1), argues that the maximum of these high-probability upper bounds on individual errors is, in fact, a high-probability upper bound on the maximum error. Thus, we can tackle error estimation without ever needing any knowledge of or bounds on the size/complexity of the input class.
The rest of the paper is organized as follows: Section 1.1 provides a comprehensive review of related literature. Section 2 presents the core theoretical framework for error estimation, structured into three key subsections.
Section 2.1 introduces the fundamental setting and notation. Section 2.2 presents the main idea in our approach. Building upon this foundation, Section 2.3 proposes an advanced localization procedure to refine the bounds derived in Section 2.2.
Section 3 then demonstrates how to apply the error estimation method in practice. We first outline a general recipe in Section 3.1, followed by two examples. In Section 3.2, we apply the error estimation theory derived in Section 2 to bound the maximum error in the multiple mean setting. Section 3.3 illustrates how to extend the theory to bound excess risk in supervised learning settings. We conclude by discussing the broader implications of our work.
Additionally, in Appendix C we provide an extension of the error estimation model to multiple hypothesis testing. Furthermore, Appendix D presents a key application of our work to contextual bandits, a challenging problem in adaptive experimentation. Contextual bandit algorithms require a measure of estimation error/uncertainty to guide exploration. While existing algorithms (with tight theoretical guarantees) rely on mathematically derived class complexity bounds to quantify such errors (Foster & Rakhlin, 2023), our error estimation approach can achieve tight error/uncertainty bounds without needing such a priori mathematical analysis. [^2] For many classes of interest, such as neural networks, we don’t have a precise mathematical characterization of class complexity and this is still an active area of research. Building on insights from (Foster et al., 2020; Krishnamurthy et al., 2023, 2024), we show how error estimation enables the development of theoretically sound contextual bandit algorithms that accommodate a wide range of policy classes.
1.1 Related Work
Error estimation is closely tied to generalization bounds in statistical learning theory. These bounds quantify the gap between a model’s empirical performance on training data and its expected performance on unseen data, uniformly across all models in a specified class. Such bounds typically rely on measures or bounds on the complexity of the model class (Shalev-Shwartz & Ben-David, 2014; Koltchinskii, 2011).
Bootstrapping is a popular empirical alternative for constructing confidence intervals, particularly effective in the case of large but finite classes, such as when estimating the mean of a high-dimensional random variable. When statistical complexity is bounded by a polynomial function of sample size, bootstrap-based methods have been applied to construct asymptotically valid confidence intervals around the maximum ($L^{\infty}$) error (see Giné & Zinn, 1990; Chernozhukov et al., 2013, 2023). However, bootstrapping faces several limitations: (1) it is effective in high but not infinite dimensions, limiting its applicability in continuous class settings, (2) it only provides asymptotic guarantees, and (3) the number of bootstrap samples, $B$ must be chosen heuristically and scales poorly with model complexity. This becomes computationally expensive as bootstrapping requires solving $B$ optimization problems. Our approach, on the other hand, simplifies to solving a single optimization problem.
The literature on PAC-Bayes bounds (see Alquier, 2021, for a user friendly tutorial) is also motivationally related to error estimation. PAC-Bayes seeks to improve upon class complexity-based bounds in traditional statistical learning theory by leveraging a prior over the given hypothesis class. With an informed prior that has a high-probability of sampling hypotheses with low average loss, the resulting generalization bounds can be much better than traditional class complexity-based bounds. However, when the prior is uninformative (e.g. uniform-like distribution over hypotheses), PAC-Bayes bounds have no advantage over traditional class complexity-based bounds. Thus, two key disadvantages with PAC-Bayes bounds are (1) the need to work with a distribution over hypotheses (ERM only outputs a single hypothesis and does not fall into this framework) and (2) a need to specify good priors. Our solution to error estimation simultaneously bounds every error of interest, and doesn’t require an informative prior to achieve an improvement over traditional methods.
The statistical learning theory bounds described above are critical for theoretically sound algorithms for active learning, contextual bandits, and reinforcement learning (RL) (Agarwal et al., 2019; Foster & Rakhlin, 2023). For contextual bandits, we replace the need to know class complexity-based bounds with a data-driven estimation of the error. One can imagine that error estimation can similarly improve on the statistical learning theory bounds in active learning and RL. Other potential areas of applications include: detecting model degradation from tracking a class of evaluation tasks, multiple hypothesis testing in genetic analysis, etc.
It is helpful to note that our work isn’t closely related to literature on constructing anytime-valid confidence intervals (CIs) (e.g., Ramdas et al., 2023). While both error estimation and the literature on anytime-valid CIs involve constructing simultaneously valid CIs, we do not explicitly model or consider time dependent structure in our data.
2 Error Estimation Theory
2.1 Preliminaries for Error Estimation
We begin by formally defining the error estimation problem. Consider an initial random variable $S_{\text{def}}$ representing a dataset. This will be the “defining” dataset, used to construct estimates and define errors. Now we introduce a set or class of indices $\mathcal{H}$ (the set of subgroups in the educational assessment example), where each element $h\in\mathcal{H}$ is an index to an estimation task with three associated quantities: an unknown target estimand $\theta_{h}$ (the true mean test scores for subgroup $h$), a given estimate $\hat{\theta}_{h}$ estimated from $S_{\text{def}}$ (the sample mean test scores), and an unknown real-valued error $e_{h}$. Here the error ($e_{h}$) is an unknown real-valued measure of “proximity” (often a normalized distance) between the given estimate ($\hat{\theta}_{h}$) and the unknown estimand ($\theta_{h}$). Let $\mathcal{E}$ be a function that given a realization $s$ of $S_{\text{def}}$, produces the class of estimand-estimate-error tuples associated with each estimation task/index. $\mathcal{E}(s):=\{(h,\hat{\theta}_{h}(s),\theta_{h},e_{h}(s))|h\in\mathcal{H}\}$, where $\hat{\theta}_{h}(s)$ is the estimate for task $h$ computed from dataset realization $s$, and $e_{h}(s)$ is the corresponding error. [^3] In fact, all quantities in $\mathcal{E}$ can depend on s, see Section 3.3, Appendix C, and Section D.2 for examples where $\mathcal{H}$ and/or $\theta_{h}$ depend on $s$ as well. To lighten the notation, we refer to $e_{h}(s)$ and $\hat{\theta}_{h}(s)$ as simply $e_{h}$ and $\hat{\theta}_{h}$ where the dependence on $s$ is clear from context.
Now consider another random variable $S_{\text{err}}$, representing the “error estimation dataset”. This dataset is drawn from some distribution $\mathcal{D}_{s}$ and will be used to construct bounds on the maximum error $\max_{h\in\mathcal{H}}e_{h}(s)$. Here, $\mathcal{D}_{s}$ may be parameterized by quantities depending on the realization $s$ of $S_{\text{def}}$. [^4] For example, in reinforcement learning, future sampling distributions may depend on earlier estimates. However, in other estimation settings, estimating some parameters of a given distribution would not affect the distribution of future/holdout samples. Once $\mathcal{D}_{s}$ is defined, $S_{\text{err}}$ and its realizations do not further depend on $s$. We often enforce such conditions by ensuring that the random dataset $S_{\text{err}}$ is either a newly sampled dataset from $\mathcal{D}_{s}$ or a holdout dataset that wasn’t used to construct any of the estimates $\hat{\theta}_{h}(s)$ in $\mathcal{E}(s)$.
For any given confidence parameter $\delta\in(0,1)$, the goal of error estimation is thus to use the random dataset $S_{\text{err}}$ to estimate a high-probability (holds with probability at least $1-\delta$) upper bound on the maximum error ($\max_{h\in\mathcal{H}}e_{h}$).
Like the union bound, our solution to the error estimation problem assumes access to an estimator ($\hat{u}:(S_{\text{err}},h,\lambda)\to\mathbb{R}$) that outputs, for any single $h\in\mathcal{H}$, an upper bound on the error $e_{h}$ with probability at least $1-\lambda$ (under the distribution of $S_{\text{err}}$) (see 2.1). Our main result (Theorem 2.2) then leverages $\hat{u}$ in order to construct a high-probability upper bound on the maximum error (that is, $\max_{h\in\mathcal{H}}e_{h}$) – without needing bounds on the complexity or size of the class $\mathcal{H}$ as an input.
Assumption 2.1 (Upper Bounds for Individual Errors).
We have access to error upper bound estimators $\hat{u}:(S_{\text{err}}\times\mathcal{H}\times(0,1))\to\mathbb{R}$ that satisfy the following condition:
$\Pr_{S_{\text{err}}\sim\mathcal{D}}(\hat{u}(S_{\text{err}},h,\lambda)\geq e_{h}|\mathcal{E}(s))\geq 1-\lambda$
for all $h\in\mathcal{H}$ and $\lambda\in(0,1)$.
2.2 Main Result: Data-Driven Error Estimation
We now describe our surprisingly simple and hence very general solution (Theorem 2.2) to the error estimation problem, which notably doesn’t require any explicit information on the size/complexity of the class $\mathcal{H}$ [^5] The validity of our solution doesn’t require any condition on the class $\mathcal{H}$ to hold. For instance, it accommodates any measure of complexity of $\mathcal{H}$ to be infinitely large.
Theorem 2.2 (Data-Driven Upper Bound on Max Error).
Suppose 2.1 holds. Let $\hat{\xi}$ denote our estimated upper bound on the maximum error, defined in (2). Then for any $\delta\in(0,1)$, we have that (3) and (4) hold. [^6] Neither (2) (definition of $\hat{\xi}$) or (3) (main guarantee) explicitly depend on the complexity size of class $\mathcal{H}$.
$\hat{\xi}(S_{\text{err}},\delta):=\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$ (2)
$\Pr_{S_{\text{err}}\sim\mathcal{D}}\Big{(}\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}\Big{|}\mathcal{E}(s)\Big{)}\geq 1-\delta$ (3)
$\Pr_{(S_{\text{def}},S_{\text{err}})}\Big{(}\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}\Big{)}\geq 1-\delta$ (4)
Proof.
Consider an index achieving maximum error, $\tilde{h}\in\arg\max_{h\in\mathcal{H}}e_{h}$ and event $\mathcal{W}(S):=\{\hat{u}(S,\tilde{h},\delta)\geq e_{\tilde{h}}\}$. We first show that $\mathcal{W}(S_{\text{err}})$ implies that $\hat{\xi}(S_{\text{err}},\delta)$ upper bounds the maximum error.
$\mathcal{W}(S_{\text{err}})$ $\implies e_{\tilde{h}}\stackrel{{\scriptstyle(i)}}{{\leq}}\hat{u}(S_{\text{err}},\tilde{h},\delta)\stackrel{{\scriptstyle(ii)}}{{\leq}}\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$ (5)
$\stackrel{{\scriptstyle(iii)}}{{\implies}}\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}.$
Here (i) follows from $\mathcal{W}(S_{\text{err}})$, (ii) follows from $\tilde{h}\in\mathcal{H}$, and (iii) follows from (2) (definition of $\hat{\xi}$) and definition of $\tilde{h}$. We now utilize (5) and 2.1 to show that (3) holds.
$\Pr_{S{\text{err}}\sim\mathcal{D}}\Big{(}\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}\Big{|}\mathcal{E}(s)\Big{)}\stackrel{{\scriptstyle(i)}}{{\geq}}\Pr_{S{\text{err}}\sim\mathcal{D}}(\mathcal{W}(S_{\text{err}})|\mathcal{E}(s))$ (6)
=$\stackrel{{\scriptstyle(ii)}}{{=}}\Pr_{S_{\text{err}}\sim\mathcal{D}}\Big{(}\hat{u}(S_{\text{err}},\tilde{h},\delta)\geq e_{\tilde{h}}\Big{|}\mathcal{E}(s)\Big{)}\stackrel{{\scriptstyle(iii)}}{{\geq}}1-\delta$
Here (i) follows from (5), (ii) follows from the definition of $\mathcal{W}$, and (iii) follows from 2.1 – showing that (3) holds. Now by the law of total expectation [^7] If we let $I=\mathbf{1}_{\{\hat{\xi}(S_{\text{err}},\delta)\geq\max_{h\in\mathcal{H}}e_{h}\}}$, the law of total expectation (or tower rule) states that $\mathbb{E}_{S_{\text{def}}}[\mathbb{E}_{S_{\text{err}}}[I|\mathcal{E}(s)]]=\mathbb{E}_{S_{\text{def}},S_{\text{err}}}[I]$ and (3), we also have that (4) holds. This completes the proof of Theorem 2.2.
Theorem 2.2 shows that, with probability at least $1-\delta$, $\hat{\xi}(S_{\text{err}},\delta)$ is an upper bound on max error ($\max_{h\in\mathcal{H}}e_{h}$). In contrast to a union bound approach, the definition of $\hat{\xi}(S_{\text{err}},\delta)$ does not explicitly depend on the complexity or size of class $\mathcal{H}$. This allows our method to naturally adapt to the correlation structure of the errors. As a consequence, our bound matches the union bound (up to constant factors) when it is tight, while providing much tighter bounds when it is loose (e.g., for highly correlated errors, see Remark 3.4 for a concrete example). Note that the dependence on $\mathcal{H}$ occurs in the maximization over the class $\mathcal{H}$. For finite classes, this maximization is straightforward, while for infinite classes it can be accomplished using standard optimization procedures such as gradient descent.
2.3 Localizing error bounds
We have introduced error estimation as a data-driven alternative to traditional bounds related to class complexity or size, such as Talagrand’s inequality (Koltchinskii, 2011; Xu & Zeevi, 2020). These methods are frequently enhanced by employing a localization argument to refine the bounds (see, for instance, (Koltchinskii, 2011)). In this section, we apply a similar localization argument to our context to achieve sharper bounds. Suppose that rather than seeking to establish an upper bound on $\max_{h}e_{h}$, we are interested in establishing a bound for the error $e_{h^{*}}$ for the instance $h^{*}:=\operatorname*{arg\,max}_{h}\theta_{h}$  [^8] In fact, $h^{*}$ can be any well-defined instance and possibly depending on $\mathcal{E}$, but independent of $S_{\text{err}}$. For simplicity, we focus on this specific case of $h^{*}:=\operatorname*{arg\,max}_{h}\theta_{h}$ as it illustrates the key idea and is sufficient for our purposes. Extensions to other choices of $h^{*}$ follow similar reasoning. In this case, rather than having to maximize an upper bound across the whole class $\mathcal{H}$, it is enough to maximize across any class that contains $h^{*}$.
To construct such classes, we introduce 2.3 which posits the existence of a lower bound on $\theta_{h^{*}}$. This assumption arises naturally in settings where we have some problem-specific characteristics which provide prior knowledge about the instance with the maximum $\theta_{h}$ value. [^9] For instance, say we are interested in upper bounding the excess risk of a given model in a class (primary goal in Section 3.3). Excess risk is defined as the maximum difference between the risk of the given model and the risk of any other model in the class – hence fits the localization setup. The maximum difference is at least zero and will be positive if the given model is not the best in class – hence satisfies 2.3 with $c=0$. . We will explore concrete examples of this in Section 3.3 and Appendix D.
Assumption 2.3 (Lower Bound for Given Instance).
We have an a priori lower bound on the maximum estimand. Specifically, there exists some constant $c$ such that $c\leq\max_{h}\theta_{h}=\theta_{h^{*}}$.
In this setting, we provide a more general alternative to Theorem 2.2, relying on a weaker alternative to 2.1 (2.4).
Assumption 2.4 (Error Estimation for Specific Instance).
We have access to error upper bound estimators $\hat{u}:(S_{\text{err}}\times\mathcal{H}\times(0,1))\to\mathbb{R}$ which satisfy
$\Pr_{S_{\text{err}}\sim\mathcal{D}}(e_{h^{*}}\leq\hat{u}(S_{\text{err}},h^{*},\delta)|\mathcal{E}(s))\geq 1-\delta$
for $h^{*}=\operatorname*{arg\,max}_{h}\theta_{h}$, and any $\delta\in(0,1)$.
Note that while 2.1 relies on the probabilistic bound $\hat{u}(S_{\text{err}},h,\delta)$ to hold for each $h\in\mathcal{H}$, 2.4 only requires a specific probabilistic bound for the instance $h^{*}$. However, since $h^{*}$ is defined in terms of unknown estimands $\theta_{h}$, we cannot directly compute $\hat{u}(S_{\text{err}},h^{*},\delta)$. Instead, our strategy is to use 2.3 to construct a non-increasing sequence of classes $\{\mathcal{H}_{k}(S_{\text{err}})\}_{k\in\mathbb{N}}$ which will all contain $h^{*}$ under certain conditions.
Corollary 2.5.
Fix $h^{*}\in\mathcal{H}$ and suppose 2.4 and 2.3 hold for some constant $c$. Further, let $\mathcal{H}_{0}:=\mathcal{H}$ and define $\mathcal{H}_{k}$, $\hat{\xi}_{k}$ inductively as in (7).
$\hat{\xi}_{k}(S_{\text{err}},\delta)$ $:=\max_{h\in\mathcal{H}_{k}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)$ (7)
$\mathcal{H}_{k+1}(S_{\text{err}})$ :=$:=\left\{h\in\mathcal{H}|-\hat{\theta}_{h}\leq\hat{\xi}_{k}(S_{\text{err}},\delta)-c\right\}$
Then, for any $\delta\in(0,1)$, we have (8) holds.
$\Pr_{(S_{\text{def}},S_{\text{err}})}\Big{(}\hat{\xi}_{k}(S_{\text{err}},\delta)\geq e_{h^{*}}\quad\forall k\in\mathbb{N}\Big{)}\geq 1-\delta$ (8)
The proof of Corollary 2.5 follows arguments similar to the proof of Theorem 2.2. See Appendix A for details.
3 Error Estimation Examples
The goal of this section is to better understand error estimation by outlining a simplified recipe for applying error estimation, followed with two examples: (i) bounding error for multiple mean estimates in Section 3.2, and (ii) bounding excess risk in Section 3.3 – which is especially useful when there isn’t a good bound on class complexity.
3.1 Simplifed Recipe for Error Estimation
Let $S_{\text{def}}$ be the dataset used for generating our estimand-estimate-error tuples and $S_{\text{err}}$ be the dataset used for producing upper bounds. In this exposition, we assume for simplicity that the errors we are interested in upper bounding are defined as $e_{h}=\theta_{h}-\hat{\theta}_{h}$, and detail how to construct explicit bounds $\hat{u}(S_{\text{err}},h,\delta)$ to satisfy 2.1.
Notation Explanation
$\theta_{h}$ Estimand for hypothesis $h$
$\hat{\theta}_{h}$ Estimate of $\theta_{h}$ using $S_{\text{def}}$
$e_{h}:=\theta_{h}-\hat{\theta}_{h}$ Error from $S_{\text{def}}$
$\hat{\theta}_{\text{err},h}$ Estimate of $\theta_{h}$ using $S_{\text{err}}$
$e_{\text{err},h}:=\theta_{h}-\hat{\theta}_{\text{err},h}$ Error from $S_{\text{err}}$
Table 1: Key Quantities for Error Estimation Framework. Note that for reasonable estimates, $e_{h}$ and $e_{\text{err},h}$ will be i.i.d. random variables from a symmetric distribution.
  1. 1.
    Generate $S_{\text{err}}$ estimates. In addition to the estimand-estimate-error tuple $(\theta_{h},\hat{\theta}_{h},e_{h})$ from $S_{\text{def}}$, we use $S_{\text{err}}$ to create estimates $\hat{\theta}_{\text{err},h}$ with corresponding errors $e_{\text{err},h}$ (see Table 1). Now, for any $h\in\mathcal{H}$,
    $e_{h}=\theta_{h}-\hat{\theta}_{\text{err},h}+\hat{\theta}_{\text{err},h}-\hat{\theta}_{h}=e_{\text{err},h}+\hat{\theta}_{h}-\hat{\theta}_{\text{err},h}$ (9)
    By our construction, $e_{h}$ and $e_{\text{err},h}$ are independent, so an instance $h^{\prime}$ determined by the realization $s$ of $S_{\text{def}}$, can be treated as a single fixed instance in $\mathcal{H}$ with respect to the randomness of $S_{\text{err}}$.
  2. 2.
    Construct $\hat{u}(S_{\text{err}},h,\delta)$ to satisfy 2.1 (or 2.4) by applying a single high-probability bound to $e_{\text{err},h}$. In other words, we set
    $\hat{u}(S_{\text{err}},h,\delta):=b_{\delta,h}+\hat{\theta}_{h}-\hat{\theta}_{\text{err},h},$
    where $b_{\delta,h}$ is a $1-\delta$ probability upper bound on a single error from $S_{\text{err}}$. Now, conditioning on $s$, under the $(1-\delta)$-probable event $\{e_{\text{err},h^{\prime}}\leq b_{\delta,h^{\prime}}\}$, [^10] Whether we rely on 2.1 or 2.4 depends on whether or not our $b_{\delta,h}$ holds as a probable upper bound for all $h$ or just for $h^{*}$. we have
    $e_{h^{\prime}}\leq b_{\delta,h^{\prime}}+\hat{\theta}_{h^{\prime}}-\hat{\theta}_{\text{err},h^{\prime}}=:\hat{u}(S_{\text{err}},h^{\prime},\delta)$ (10)
    Thus we have constructed $\hat{u}(S_{\text{err}},h,\delta)$ which satisfy 2.1, and we can set the max error bound to be $\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$.
  3. 3.
    (Optional) Localize the bound. If we are in the setting described in Section 2.3, i.e. bounding $e_{h^{*}}$, where $h^{*}:=\operatorname*{arg\,max}\theta_{h}$, and we have $\theta_{h^{*}}\geq c$ (i.e. 2.3 holds for some constant $c$), we can improve our error bound by Corollary 2.5. We define $\mathcal{H}_{0}=\mathcal{H}$, and $\mathcal{H}_{k}$, $\hat{\xi}_{k}$ as in (7). This gives
    $\Pr_{S\text{err}\sim D}(\forall k\in\mathbb{N},\hat{\xi}_{k}(S_{\text{err}},\delta)\geq e_{h^{*}}|\mathcal{E}(s))\geq 1-\delta$ (11)
    Note that although Corollary 2.5 provides a bound for any $k\in\mathbb{N}$, at some point these upper bounds will converge in practice as $k\to\infty$.
Now under the distribution of $(S_{\text{def}},S_{\text{err}})$, we make a remark tying the distribution of our constructed $\hat{u}(S_{\text{err}},h,\delta)$ (see (10)) to the distribution of $e_{h}$.
Remark 3.1 ($\hat{u}(S_{\text{err}},h,\delta)$ is often a surrogate for $e_{h}$).
Note that in this setting:
$\hat{u}(S_{\text{err}},h,\delta)=b_{\delta,h}+\hat{\theta}_{h}-\hat{\theta}_{\text{err},h}=b_{\delta,h}+e_{\text{err},h}-e_{h},$
Thus, as $e$ and $e_{\text{err}}$ are defined to behave as i.i.d. random variables, our constructed $\hat{u}$ inherits both the distributional and correlational structure of $e$ up to multiplicative and additive constant factors. [^11] If additionally $e$ is symmetrically distributed around the zero vector, then $\hat{u}$ will follow a scaled version of the same distribution. For example, if $e_{h}\sim\mathcal{N}(0,1)$, then $\hat{u}(S_{\text{err}},h,\delta)\sim\mathcal{N}(b_{\delta,h},2)$
Moreover, if $b_{\delta,h}$ is constant across $h\in\mathcal{H}$ (as in examples below),
$\text{Cov}(\hat{u}(S_{\text{err}},h,\delta),\hat{u}(S_{\text{err}},h^{\prime},\delta))=2\text{Cov}(e_{h},e_{h^{\prime}}).$
This indicates that $\hat{u}(S_{\text{err}},h,\delta)$ is a surrogate for $e_{h}$, thus using $\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$ as a bound on $\max_{h\in\mathcal{H}}e_{h}$ is often tight up to constant factors.
Remark 3.2 (There is Room for Improvement).
In an ideal world, $\hat{u}(S_{\text{err}},h,\delta)$ would be identically distributed as $e_{h}$ (while still being a valid upper bound). Our construction for $\hat{u}(S_{\text{err}},h,\delta)$ often leads to twice as much variance as $e_{h}$ and we need twice as much data to define the same error. Hence, compared to the “ideal” scenario, our error estimation intervals are larger by a factor of $\sqrt{4}=2$.
3.2 Confidence Interval Construction for Multiple Mean Estimates
Consider a random mapping $Y:\mathcal{H}\rightarrow[-M,M]$ (such as subgroups to average test score), and let $\mathcal{D}$ denote its distribution. Our target estimands are given by $\theta_{h}:=\mathop{\mathbb{E}}_{Y\sim\mathcal{D}}[Y(h)]$ for all $h\in\mathcal{H}$, and our goal is to construct CIs for these estimands that simultaneously hold with high-probability. [^12] Refer to Appendix B for a formal conversion of the subgroup example presented in Section 1 to the setting discussed in this section.We first produce estimates and define errors for each $h\in\mathcal{H}$.
[…]
Term Definition Term Definition
$\hat{\theta}_{h}$ $\sum_{y\in S_{\text{def}}}\frac{y(h)}{n}$ $\hat{\theta}_{\text{err},h}$ $\sum_{y\in S_{\text{err}}}\frac{y(h)}{n}$
$\hat{\sigma}_{h}^{2}$ $\sum_{y\in S_{\text{def}}}\frac{(y(h)-\hat{\theta}_{h})^{2}}{n-1}$ $\hat{\sigma}_{\text{err},h}^{2}$ $\sum_{y\in S_{\text{err}}}\frac{(y(h)-\hat{\theta}_{\text{err},h})^{2}}{n-1}$
$e_{h}$ $\frac{\sqrt{n}|\hat{\theta}_{h}-\theta_{h}|}{\hat{\sigma}_{h}}$ $e_{\text{err},h}$ $\frac{\sqrt{n}|\hat{\theta}_{\text{err},h}-\theta_{h}|}{\hat{\sigma}_{\text{err},h}}$
Table 2: Error-Estimation Quantities for Multiple Means
Thus, for any $h\in\mathcal{H}$, $\hat{\theta}_{h}$ is our mean estimate around which we will construct our CI. The error $e_{h}$ is then defined as the absolute difference between this estimate and the target estimand ($|\theta_{h}-\hat{\theta}_{h}|$) divided by the sample standard error ($\hat{\sigma}_{h}/\sqrt{n}$). Note that upper bounding the error $e_{h}$ gives us an interval for the estimand $\theta_{h}$. Hence, by constructing a high-probability upper bound on the max error ($\max_{h\in\mathcal{H}}e_{h}$), we get simultaneously valid high-probability CIs for all our estimands. The fundamental quantities we establish from dataset $S_{\text{err}}$ are a mean estimate $\hat{\theta}_{\text{err},h}$, sample standard deviation $\hat{\sigma}_{\text{err},h}$, and $e_{\text{err},h}$ (see Table 2). We use these quantities when applying Theorem 2.2 for error estimation.
We now derive an estimate $\hat{u}$ satisfying 2.1. That is, for any $h\in\mathcal{H}$, $\hat{u}(S_{\text{err}},h,\delta)$ should be an upper bound on the input error $e_{h}$ with probability at least $1-\delta$. In this section, we rely on the central limit theorem (CLT) to construct $\hat{u}$ that satisfies this guarantee asymptotically (a sample size of $n\geq 30$ is often considered sufficient for the asymptotic guarantees of CLT to hold). [^13] Our choice to rely on CLT here is due to its popularity in practice. To get non-asymptotic results, one may construct $\hat{u}$ based on Hoeffding or Bernstein’s inequality. By construction, we have defined $e_{h}$ and $e_{\text{err},h}$ to be distributed as $|\mathcal{N}(0,1)|$ random variables. Thus, from the CLT, $z_{\delta/2}$ upper bounds $e_{\text{err},h}$ asymptotically with probability at least $1-\delta$, where $z_{\delta}$ is the $\delta$ critical value for the standard normal distribution ($N(0,1)$). [^14] That is, a sample from the standard normal distribution is greater than $z_{\delta}$ with probability $\delta$. Now,
$e_{\text{err},h}\leq z_{\delta/2}$ (12)
$\implies$ $e_{h}=\frac{\sqrt{n}|\hat{\theta}_{h}-\theta_{h}|}{\hat{\sigma}_{h}}$
$\leq\frac{\sqrt{n}|\hat{\theta}_{h}-\hat{\theta}_{\text{err},h}|}{\hat{\sigma}_{h}}+\frac{\hat{\sigma}_{\text{err},h}z_{\delta/2}}{\hat{\sigma}_{h}}=:\hat{u}(S_{\text{err}},h,\delta)$
Hence, for all $h$, $\Pr_{S_{\text{err}}}(\hat{u}(S_{\text{err}},h,\delta)\geq e_{h}|S_{\text{def}})\geq 1-\delta$ asymptotically holds. Thus when $n$ is large enough for CLT-based CIs to hold, we have 2.1 holds. Hence from Theorem 2.2, we have the following corollary.
Corollary 3.3.
Consider the setup in Section 3.2. Suppose the following CLT-based CIs hold: $\Pr_{S_{\text{err}}}(\hat{u}(S_{\text{err}},h,\delta)\geq e_{h}|S_{\text{def}})\geq 1-\delta$ for all $h\in\mathcal{H}$. Then for $\hat{\xi}(S_{\text{err}},\delta):=\max_{h\in\mathcal{H}}\hat{u}(S_{\text{err}},h,\delta)$, we have (13) holds.
$\Pr_{\bar{S}}\left(\forall h\in\mathcal{H},\;|\hat{\theta}_{h}-\theta_{h}|\leq\frac{\hat{\sigma}_{h}}{\sqrt{n}}\hat{\xi}(S_{\text{err}},\delta)\right)\geq 1-\delta$ (13)
Proof.
From Theorem 2.2, we have $\Pr_{\bar{S}}\left(\max_{h\in\mathcal{H}}e_{h}\leq\hat{\xi}(S_{\text{err}},\delta)\right)\geq 1-\delta$. The result follows from noting that $\max_{h\in\mathcal{H}}e_{h}\leq\hat{\xi}(S_{\text{err}},\delta)$ is equivalent to $\forall h\in\mathcal{H},\;|\hat{\theta}_{h}-\theta_{h}|\leq\frac{\hat{\sigma}_{h}}{\sqrt{n}}\hat{\xi}(S_{\text{err}},\delta)$. ∎
We are able to construct simultaneously valid CIs without requiring explicit union-bound-based adjustments. Using $\hat{\xi}(S_{\text{err}},\delta)$ in our CIs instead of $z_{\delta/2}$ is a data-driven adjustment for constructing multiple simultaneously valid CIs. This can be much better than the union-bound-based correction of $z_{\delta/(2|\mathcal{H}|)}$ when $\mathcal{H}$ is finite. To see the benefits of the above data-driven adjustment, see Figure 1.
Remark 3.4 (Tightness compared to union bound).
Our error estimation bound is the maximum of $\hat{u}(S_{\text{err}},h,\delta)$, to further bound this analytically we can use the union bound. Under the randomness of $\bar{S}$, one can show that for the multiple mean setting $\hat{u}(S_{\text{err}},h,\delta)$ is asymptotically distributed as $z_{\delta/2}+|N(0,2)|$. Applying the union bound, we get our error estimation bound is further bounded by $O(\sqrt{\log|\mathcal{H}|})$ – showing that, up to constant factors, error estimation is no worse than applying a direct union bound on errors. [^15] For the multiple mean setting, $e_{h}$ is asymptotically distributed as $|N(0,1)|$. Further, these bounds are tight (up to constants) when our mean estimation tasks in $\mathcal{H}$ are independent. Thus, although our approach doesn’t rely on $|\mathcal{H}|$ as an input, our bound may still grow as a function of $|\mathcal{H}|$. In practice, our estimation tasks are not independent, and in such cases our method can provide significantly tighter bounds by escaping the explicit dependence on complexity or size of $\mathcal{H}$.
Before moving on to the infinite class setting, we visualize an example of this tightness in Figure 1 where we errors are correlated, and we see that as $|\mathcal{H}|$ grows, the union bound becomes progressively looser, while our error bound remains stable. We also see that for smaller group sizes $|\mathcal{H}|$ our bound may be looser, as we pay a constant factor price (see Remark 3.2).
Figure 1: In this plot, we bound maximum error across estimates quantile estimates. We use $n=10000$ i.i.d. samples from a standard normal distribution to estimate an increasing number of quantiles. We use two methods to bound the maximum error, and plot the excess (bound - true maximum error) of error bounds vs. increasing number of quantiles estimated for each: traditional union bound (orange) and our error estimation method (blue).
3.3 Supervised Learning: Excess Risk Estimation and Constructing Set of Potentially Optimal Models
We consider the supervised learning setting – described by a distribution $\mathcal{D}$ over $\mathcal{X}\times\mathcal{Y}$, a model class $\mathcal{G}\subseteq\mathcal{X}\rightarrow\mathcal{Y}$, a loss function $l:\mathcal{X}\times\mathcal{Y}\times\mathcal{G}\rightarrow[0,M]$ (for some $M>0$), and two independent datasets $S_{\text{def}}$ and $S_{\text{err}}$ each containing $n$ i.i.d. samples from the distribution $\mathcal{D}$. For any model $g$, the risk is defined as $\mathcal{R}(g):=\mathop{\mathbb{E}}_{D}[l(X,Y,g)]$. The optimal model $g^{*}\in\mathcal{G}$ is then a solution to $\min_{g\in\mathcal{G}}\mathcal{R}(g)$ and the excess risk is defined as $\Delta(g):=\mathcal{R}(g)-\mathcal{R}(g^{*})$. Excess risk ($\Delta(g)$) can be interpreted as a measure of room for improvement over the model $g$.
Suppose we are motivated by the need to determine if we should collect more data for learning from the class $\mathcal{G}$. In this case, for some estimated model $g_{\text{def}}\in\mathcal{G}$ (which may depend on $S_{\text{def}}$ but not on $S_{\text{err}}$), we would like to estimate an upper bound on excess risk ($\Delta(g_{\text{def}})$). If the upper bound on excess risk is large, then collecting more data should allow us to learn a better model from the class $\mathcal{G}$, or at least help us decrease uncertainty for model quality. In general, bounding excess risk is a challenging problem, with many existing approaches relying on naive methods or complex theoretical tools. Using the localization technique introduced in Section 2.3, we construct excess risk bounds and are thus also be able to construct a set of potentially optimal models – that is, a set which will contain $g^{*}$ with high-probability. [^16] Constructing a set of potentially optimal models is the primary statistical step when constructing confidence intervals for UCB (upper confidence bound) based contextual bandit/RL algorithms.
We formulate the problem of upper bounding the excess risk ($\Delta(g_{\text{def}})$) as an error estimation problem. For each $g\in\mathcal{G}$, we define our estimand as $\theta_{g}:=\mathcal{R}(g_{\text{def}})-\mathcal{R}(g)=\mathop{\mathbb{E}}_{D}[l(X,Y,g_{\text{def}})-l(X,Y,g)]$. That is, our estimand $\theta_{g}$ measures the sub-optimality of $g_{\text{def}}$, when compared against a candidate model $g$. Therefore, the excess risk $\Delta(g_{\text{def}})$ is equal to $\theta_{g^{*}}$. Now, our estimate is defined as $\hat{\theta}_{g}:=(1/n)\sum_{(x,y)\in S_{\text{def}}}(l(x,y,g_{\text{def}})-l(x,y,g))$ and the corresponding error is $e_{g}:=\theta_{g}-\hat{\theta}_{g}$. We define a corresponding estimate from $S_{\text{err}}$, $\hat{\theta}_{g,\text{err}}$ as outlined in Table 3.
Term Definition
$\theta_{g}$ $\mathcal{R}(g_{\text{def}})-\mathcal{R}(g)=\mathop{\mathbb{E}}_{D}[l(X,Y,g_{\text{def}})-l(X,Y,g)]$
$\hat{\theta}_{g}$ $(1/n)\sum_{(x,y)\in S_{\text{def}}}(l(x,y,g_{\text{def}})-l(x,y,g))$
$\hat{\theta}_{g,\text{err}}$ $(1/n)\sum_{(x,y)\in S_{\text{err}}}(l(x,y,g_{\text{def}})-l(x,y,g))$
$e_{g}$ $\theta_{g}-\hat{\theta}_{g}$
Table 3: Error Estimation Quantities for Excess Risk Estimation
We now construct $\hat{u}$. From Hoeffding’s inequality, we have
$\hat{u}(S_{\text{err}},g,\delta):=\hat{\theta}_{g,\text{err}}-\hat{\theta}_{g}+M\sqrt{\frac{\log(1/\delta)}{2n}}$
satisfies 2.1. Now, defining $\hat{\xi}_{0}(S_{\text{err}},\delta):=\max_{g\in\mathcal{G}}\hat{u}(S_{\text{err}},g,\delta)$, we have by Theorem 2.2, (14) holds with probability at least $1-\delta$.
$\Delta(g_{\text{def}})=\theta_{g^{*}}=\hat{\theta}_{g^{*}}+e_{g^{*}}\leq\hat{\theta}_{g^{*}}+\hat{u}(S_{\text{err}},g^{*},\delta)$ (14)
$\leq\max_{g\in\mathcal{G}}\hat{\theta}_{g}+\hat{\xi}_{0}(S_{\text{err}},\delta)$
By the definition of $g^{*}$, $\theta_{g^{*}}=\max_{g\in\mathcal{G}}\theta_{g}$, and $\theta_{g^{*}}(=\mathcal{R}(g_{\text{def}})-\mathcal{R}(g^{*}))\geq 0$, so we have satisfied 2.3 with $c=0$. We can thus apply Corollary 2.5 to improve the bound. Let $\mathcal{G}_{0}=\mathcal{G}$, and define
$\hat{\xi}_{k}(S_{\text{err}},\delta)$ $:=\max_{g\in\mathcal{G}_{k}}\hat{u}(S_{\text{err}},g,\delta);$ (15)
$\mathcal{G}_{k}$ :=$:=\bigg{\{}g\in\mathcal{G}_{k-1}\bigg{|}-\hat{\theta}_{g,\text{def}}\leq\hat{\xi}_{k-1}(S_{\text{err}},\delta)\bigg{\}};$
For any $x\in\mathcal{X}$, this set of potentially optimal models ($\mathcal{G}_{k}$) allows us to construct a high-probability confidence set for $g^{*}(x)$, given by $\{g(x)|g\in\cap_{k\in\mathbb{N}}\mathcal{G}_{k}\}$. [^17] For some $K$, one would observe $\hat{\xi}_{K+1}=\hat{\xi}_{K}$, and stop the localization procedure. Such confidence sets can be helpful for a variety of decision-making tasks. We summarize our results in Corollary 3.5. It is important to note that unlike popular theoretical bounds on excess risk (e.g., Koltchinskii, 2011), none of our results required the complexity of class $\mathcal{G}$ as an input.
Corollary 3.5.
Consider the supervised learning setting described in Section 3.3. With probability at least $1-\delta$, we have (16) hold for all $k\in\mathbb{N}$.
$\forall g\in\mathcal{G},\;\;e_{g}\leq\hat{\xi}_{0}(S_{\text{err}},\delta);\;\;\;\;\;g^{*}\in\mathcal{G}_{k};\;\;\;$ (16)
$\forall x\in\mathcal{X},\;\;g^{*}(x)\in\{g(x)|g\in\mathcal{G}_{k}\};$
$\Delta(g_{\text{def}})\leq\max_{g\in\mathcal{G}_{k}}\hat{\theta}_{g}+\hat{\xi}_{k}(S_{\text{err}},\delta).$
Remark 3.6.
Recall $\hat{\theta}_{g}:=(1/n)\sum_{(x,y)\in S_{\text{def}}}(l(x,y,g_{\text{def}})-l(x,y,g))$, and so a reasonable choice of $g_{\text{def}}$ would be the minimizer $g_{\text{def}}=\operatorname*{arg\,min}_{g\in\mathcal{G}}\mathbb{E}_{S_{\text{def}}}[l(X,Y,g)]$. In this case we would have $\hat{\theta}_{g}\leq 0$ and we could reduce our excess risk bound to $\Delta(g_{\text{def}})\leq\min_{k}\hat{\xi}_{k}(S_{\text{err}},\delta)$.
Figure 2 illustrates the performance of our error estimation bound on excess risk compared to the empirical risk of $g_{\text{def}}$ on $S_{\text{err}}$ for increasing noise levels. Empirical risk is an estimate of true risk, which serves as valid yet naive bound on the excess risk. This is because the true risk is the sum of the excess risk and the irreducible noise of the dataset. Our error estimation approach is designed to cancel out irreducible noise via the construction $\hat{u}$, as the performance of $g_{\text{def}}$ is compared to other models in $\mathcal{G}$. As the noise levels increase, our (probabilistic) bound shows increased variance, but remains closer on average to the true excess risk, demonstrating its ability to distinguish between excess risk and irreducible noise. See Section B.1 for details regarding the simulation.
Figure 2: Comparison of excess risk bounds over a class of Neural Networks. We plot the empirical risk of $g_{\text{def}}$ on $S_{\text{err}}$ (blue), our (probabilistic) error estimation bound (orange) and the true excess risk (green) against increasing noise levels.
4 Conclusion
The problem of bounding the maximum error for a class of estimand-estimate-error tuples (i.e. error estimation problem) is fundamental to statistical learning theory and is applicable to a wide range of important problems. For large and infinite classes, the error estimation problem is either computationally expensive or requires mathematically analyzing class complexity. In this paper we circumvent such challenges by solving the error estimation problem through a single optimization procedure, without needing to mathematically analyze the complexity of the underlying class. Our procedure can further be combined with localization arguments to more tightly bound errors for specific hypotheses of interest (e.g. in supervised learning, we are particularly interested in the hypothesis that minimizes the true risk).
We show error estimation can be used to: construct multiple simultaneously valid CIs, develop multiple-hypothesis testing procedures, bound excess risk in supervised learning, construct the set of all potentially optimal models in supervised learning, construct point-wise confidence sets that contain the prediction of the optimal supervised learning model, and enable the development of a flexible and computationally efficient contextual bandit pipeline.
Error estimation is especially valuable in the contextual bandit setting, where uncertainty quantification plays a key role in the performance of the algorithm. By leveraging the flexibility of error estimation, we are able to construct a theoretically sound bandit pipeline which can use any black box regression and policy learning subroutine, and allows practitioners to use these subroutines without needing to theoretically analyze the sub-routine or the class being learned from (see Appendix D).
Appendix A Proof of Corollary 2.5
Suppose 2.4 holds, and consider the event $\mathcal{W}^{*}(S_{\text{err}})=\{e_{h^{*}}\leq\hat{u}(S_{\text{err}},h^{*},\delta)\}$, and a class $\mathcal{G}(S_{\text{err}})\subseteq\mathcal{H}$ such that $\mathcal{W}^{*}(S_{\text{err}})\subseteq\{h^{*}\in\mathcal{G}(S_{\text{err}})\}$. We then have
$\mathcal{W}^{*}(S_{\text{err}})\implies e_{h^{*}}\stackrel{{\scriptstyle(i)}}{{\leq}}\hat{u}(S_{\text{err}},h^{*},\delta)\stackrel{{\scriptstyle(ii)}}{{\leq}}\max_{h\in\mathcal{G}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)$ (17)
Here (i) follows from the definition of $\mathcal{W}^{*}(S_{\text{err}})$, and (ii) follows from the condition placed on $\mathcal{G}(S_{\text{err}})$. We now use (17) to show that
$\Pr\left(e_{h^{*}}\leq\max_{h\in\mathcal{G}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)|\mathcal{E}(s)\right)\stackrel{{\scriptstyle(i)}}{{\geq}}\Pr_{S{\text{err}}\sim\mathcal{D}}(\mathcal{W}^{*}(S_{\text{err}})|\mathcal{E}(s))\stackrel{{\scriptstyle(ii)}}{{=}}\Pr_{S{\text{err}}\sim\mathcal{D}}\Big{(}\hat{u}(S_{\text{err}},h^{*},\delta)\geq e_{h^{*}}\Big{|}\mathcal{E}(s)\Big{)}\stackrel{{\scriptstyle(iii)}}{{\geq}}1-\delta$ (18)
Here (i) holds by (17), (ii) holds by the definition of $\mathcal{W}^{*}(S_{\text{err}})$, and (iii) holds by 2.4. Furthermore, if we can construct a sequence of classes $\{\mathcal{H}_{k}(S_{\text{err}})\}_{k=0}^{\infty}\subseteq\mathcal{H}$ where for each $k$, $\mathcal{W}^{*}(S_{\text{err}})\subseteq\{h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\}$, we can consider $\mathcal{H}_{\infty}$, and apply (18) 2.4, $\Pr\left(e_{h^{*}}\leq\max_{h\in\cap_{k}\mathcal{H}_{k}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta)\right)\geq 1-\delta$.
We construct $\mathcal{H}_{k}$ as follows. Let $\mathcal{H}_{0}(S_{\text{err}})=\mathcal{H}$. Now define (as in (7))
$\hat{\xi}_{k}(S_{\text{err}},\delta):=\max_{h\in\mathcal{H}_{k}(S_{\text{err}})}\hat{u}(S_{\text{err}},h,\delta),\;\;\mathcal{H}_{k+1}(S_{\text{err}}):=\left\{h\in\mathcal{H}|-\hat{\theta}_{h}\leq\hat{\xi}_{k}(S_{\text{err}},\delta)-c\right\}$ (19)
We claim $\mathcal{W}^{*}(S_{\text{err}})\subseteq\{h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\}$ for all $k\in\mathbb{N}$. To show this we assume $\mathcal{W}^{*}(S_{\text{err}})$ holds and deduce $h^{*}\in\mathcal{H}_{k}(S_{\text{err}})\forall k$ by induction. First note that this holds true for $k=0$ since $h^{*}\in\mathcal{H}_{0}(S_{\text{err}})=\mathcal{H}$. Now, assume $h^{*}\in\mathcal{H}_{k}(S_{\text{err}})$. Recall that $e_{h^{*}}=\theta_{h^{*}}-\hat{\theta}_{h^{*}}\geq c-\hat{\theta}_{h^{*}}$ by 2.3. Thus, by definition of $\mathcal{H}_{k+1}(S_{\text{err}})$, we have $h^{*}\in\mathcal{H}_{k+1}(S_{\text{err}})$. Thus we have arrived at the following localization corollary.
Appendix B Additional details for Section 3.2
In this section, we show how to formally convert the example in Section 1 into the setting described in Section 3.2. Recall the goal of constructing simultaneously valid CIs on average test scores for subgroups (e.g.“female students”, “fourth-graders”, etc.) indexed by $\mathcal{H}$ [^18] With mild abuse of notation, we let any $h\in\mathcal{H}$ denote both the subgroup index and subgroup itself. More formally, let $(X,T)\in\mathcal{X}\times[-1,1]$ be a pair of random variables corresponding to a randomly sampled student’s feature vector and their test score respectively. The average test score for any subgroup $h$ is then given by $\theta_{h}$ in (20).
$\theta_{h}:=\mathop{\mathbb{E}}[T|X\in h]=\mathop{\mathbb{E}}[Y(h)],$ (20)
$Y(h):=\frac{T\cdot\mathbf{1}(X\in h)}{p_{h}},\;p_{h}:=\Pr(X\in h).$
For ease of exposition, we assume the probability ($p_{h}$) of a random student belonging to subgroup $h$ is known to us and is lower bounded by $1/M$ for some $M\in[1,\infty)$. [^19] We can always estimate $p_{h}$ from $S_{\text{def}}$ and reset $\mathcal{H}\leftarrow\{h\in\mathcal{H}|p_{h}\geq 1/M\}$Corollary 3.3 would still hold. The remarks would also hold if cross-fitting (Newey & Robins, 2018) were used to account for unknown $p_{h}$ in estimates from $S_{\text{def}}$ and $S_{\text{err}}$.
Although the errors in this setting, unlike in Section 3.1, are normalized, we still observe that $\hat{u}$ inherits the distribution of $e$. We formalize this in the following remark.
Remark B.1 (Asymptotic Distribution of $\hat{u}$).
As $|\bar{S}|\rightarrow\infty$, we have $\hat{u}(S_{\text{err}},h,\delta)$ converges in distribution to $|\mathcal{N}(0,2)|+z_{\delta/2}$. Note that in this setting, $e_{h}$ converges in distribution to $|\mathcal{N}(0,1)|$.
Proof.
$\hat{u}(S_{\text{err}},h,\delta)=\Bigg{|}\frac{\sqrt{n}((\hat{\theta}_{h}-\theta_{h})+(\theta_{h}-\hat{\theta}_{\text{err},h})}{\hat{\sigma}_{h}}\Bigg{|}+\frac{\hat{\sigma}_{\text{err},h}z_{\delta/2}}{\hat{\sigma}_{h}}$ (21)
$\rightarrow|\mathcal{N}(0,2)|+z_{\delta/2}.$
[…]
B.1 Details for Figures
  • Figure 1 is generated as follows:
    We generate $n=10000$ i.i.d. samples from a $\mathcal{N}(0,1)$ distribution, and empirically estimate the quantiles of this distribution. Next, we define our errors to be the absolute value of the difference between these estimates and the true theoretical quantiles, normalized so that each error is $|\mathcal{N}(0,1)|$ (to do this, we rely on the distribution of a sample quantile, $\hat{\theta}_{q}\sim\mathcal{N}\left(\theta_{q},\frac{q(1-q)}{nf(\theta_{q})^{2}}\right)$. As we don’t assume access to the density function $f$, we estimate this using a gaussian kernel estimator. Once we have defined our errors, we use the error estimation theory above, along with an additional dataset of $n=10000$ points to create bounds on the maximum error. We compare this to the bounds generated by union bounding for the original dataset. We repeat each experiment 100 times and plot the average.
  • Figure 2 is generated as follows:
    We define a model class of Neural Networks, then use the Iris dataset from UCI ML Repo as our dataset. We first train a model $g^{*}$ on the dataset to serve as the true realizable model, so we alter the data to be generated by the model. We then add noise to the dataset, split it into $S_{\text{def}}$ and $S_{\text{err}}$, and train a secondary model $g_{est}$ on $S_{\text{def}}$. Next we evaluate the risk of model $g_{\text{def}}$ over increasingly noisy versions of $S_{\text{def}}$ and $S_{\text{err}}$, and use the error estimation theory to create bounds on excess risk. To evaluate the excess risk bound, we iteratively search for a model that maximizes the difference in empirical risk under $S_{\text{def}}$ and $S_{\text{err}}$, subject to an increasingly strict localization constraint. Additionally, instead of relying on Hoeffding’s inequality we use conformal methods to estimate a high probability upper bound on a single error.
Appendix C Multiple Hypothesis Testing
In hypothesis testing, we seek to compare a real-valued estimand of interest against a specified threshold. The null hypothesis claims the estimand is equal to the threshold, and the alternative hypothesis claims the estimand is greater than the threshold. [^20] This choice of alternative hypothesis is without loss of generality. One may also consider an alternative hypothesis where the estimand is less than the threshold, or the estimand is not equal to the threshold. Given a set $\mathcal{H}$ of hypotheses, a standard goal is to reject as many null hypotheses as possible while controlling the family-wise error rate (FWER) to be at most some $\delta\in(0,1)$. [^21] FWER is the probability of incorrectly rejecting at least one null hypothesis. A popular approach to this problem is the Bonferroni correction, which can be seen as a generalization of the union bound in that it allows flexibility in assigning weights to different hypotheses. In this section, we use error-estimation to provide a data-driven alternative to the weighted Bonferroni correction.
[…]
Corollary C.1.
[…]
$\mathcal{H}_{r}:=\bigg{\{}h\in\mathcal{H}\bigg{|}\iota(h)=1,\hat{\theta}_{h}>\frac{1}{b(h)}\frac{\hat{\sigma}_{\text{def},h}}{\sqrt{n}}\hat{\xi}_{w}(S_{\text{err}},\delta)\bigg{\}},$ (22)
$\Pr_{\bar{S}}\Big{(}\forall h\in\mathcal{H}_{r},\;\theta_{h}>0\Big{)}\geq 1-\delta.$ (23)
Proof.
Following the argument in Corollary 3.3 (using Theorem 2.2), we have:
$\Pr_{\bar{S}}\Big{(}\max_{\{h\in\mathcal{H}|\iota(h)=1\}}b(h)e_{h}\leq\hat{\xi}_{w}(S_{\text{err}},\delta)\Big{)}\geq 1-\delta$ (24)
$\equiv$ $\Pr_{\bar{S}}\Big{(}\forall h\in\mathcal{H}\text{ s.t. }\iota(h)=1,|\hat{\theta}_{h}-\theta_{h}|\leq\frac{1}{b(h)}\frac{\hat{\sigma}_{\text{def},h}}{\sqrt{n}}\hat{\xi}_{w}(S_{\text{err}},\delta)\Big{)}\geq 1-\delta$
$\implies$ $\Pr_{\bar{S}}\Big{(}\forall h\in\mathcal{H}\text{ s.t. }\iota(h)=1,\theta_{h}\geq\hat{\theta}_{h}-\frac{1}{b(h)}\frac{\hat{\sigma}_{\text{def},h}}{\sqrt{n}}\hat{\xi}_{w}(S_{\text{err}},\delta)\Big{)}\geq 1-\delta$
$\implies$ $\Pr_{\bar{S}}\Big{(}\forall h\in\mathcal{H}_{r},\;\theta_{h}>0\Big{)}\geq 1-\delta.$
Note that FWER for Corollary C.1 is at most $\delta$ for any choice of $w$. Now let us briefly discuss the choices for $w=(\iota,b)$. A simple choice is to let $b\equiv\iota\equiv 1$, this choice makes Corollary C.1 a direct application of Corollary 3.3. However, while keeping FWER less than $\delta$, a more intelligent choice for $w$ will allow us to discover more $h$ for which the alternative hypothesis is true (larger $\mathcal{H}_{r}$). Note that even for rejecting the single null hypothesis $\theta_{h}=0$ with dataset $S_{\text{def}}$, we would require $\hat{\theta}_{h}>\hat{\sigma}_{\text{def},h}z_{\delta}/\sqrt{n}$. Hence for multiple testing, we only need to focus our attention on $h$ satisfying this condition. It would thus be reasonable to choose $\iota(h)=\mathbf{1}(\hat{\theta}_{h}>\hat{\sigma}_{\text{def},h}z_{\delta}/\sqrt{n})$, ensuring $\hat{\xi}_{w}(S_{\text{err}},\delta)$ is not larger than necessary. Using $b(h)=\hat{\sigma}_{\text{def},h}\min(1,1/|\hat{\theta}_{h}|)$ is also a reasonable choice – this makes the tests in (22) more comparable – allowing for $\mathcal{H}_{r}$ to be much larger when possible. If we are willing to incur a larger FWER in exchange for more nulls rejected, we can set $\iota(h)$ as zero more often and extrapolate the test in (22) to such $h$. Such approaches could allow us to tackle objectives like controlling the false discovery rate (FDR). Analyzing the choices of $w$ and its various implications is beyond the scope of this paper.
Appendix D Contextual Bandit Pipeline with Error Estimation
In this section, we describe a modular contextual bandit pipeline that leverages error estimation. We begin by introducing the standard stochastic contextual bandit setting, defined by a context space $\mathcal{X}$, a finite set of arms $\mathcal{A}=[K]$, and a fixed but unknown distribution $\mathcal{D}$ over contexts and arm rewards. [^23] Here $[n]$ denotes the set $\{1,2,\dots,n\}$. A contextual bandit algorithm (learner) interacts with samples from this unknown distribution $\mathcal{D}$ over a sequence of rounds. At each round $t$, a context $x_{t}$ and a reward vector $r_{t}\in[0,1]^{\mathcal{A}}$ are sampled from the joint distribution $\mathcal{D}$. Initially, only the context $x_{t}$ is revealed to the learner. To make a decision, the learner utilizes data from prior rounds to construct an “exploration policy” $p_{t}$ – which is a probability kernel mapping $\mathcal{X}\times\mathcal{A}$ to $[0,1]$. The learner then chooses an arm $a_{t}$ sampled from the distribution $p_{t}(\cdot|x_{t})$ and observes the corresponding reward $r_{t}(a_{t})$, this data is later used for constructing exploration policies in subsequent rounds.
Key concepts. Before defining our objectives, we introduce several key concepts. A reward model $f$ maps $\mathcal{X}\times\mathcal{A}$ to $[0,1]$, with $f^{*}(x,a):=\mathop{\mathbb{E}}_{D}[r_{t}(a)|x_{t}=x]$ denoting the true conditional expected reward model. A (deterministic) policy $\pi$ is a map from $\mathcal{X}$ to $\mathcal{A}$, and a randomized policy $p$ is a probability kernel mapping $\mathcal{X}\times\mathcal{A}$ to $[0,1]$. For a given model $f$ and a randomized policy $p$, we denote the expected instantaneous reward of $p$ with respect to model $f$ as $R_{f}(p)$ – formally defined in (25). We denote the true expected reward $R_{f^{*}}(p)$ as $R(p)$ to simplify notation when there is no risk of confusion. We also let $V(p,\pi)$ denote the “cover” of policy $\pi$ under data collected by a (randomized) policy $p$, which is formally defined in (25) – it is proportional to the variance of evaluating the policy $\pi$ using data collected by the policy $p$. Lower values of the cover indicate that data collected under $p$ provides more precise and reliable estimates for the performance of $\pi$.
$R_{f}(p):=\mathop{\mathbb{E}}_{(x,r)\sim D}\mathop{\mathbb{E}}_{a\sim p(\cdot|x)}[f(x,a)],\quad\text{and }V(p,\pi):=\mathop{\mathbb{E}}_{(x,r)\sim D,a\sim\pi(\cdot|x)}\bigg{[}\frac{\pi(a|x)}{p(a|x)}\bigg{]}.$ (25)
Our learner (bandit algorithm) is given a policy class $\Pi$. Let $\pi^{*}$ denote the optimal policy in the class $\Pi$, that is $\pi^{*}\in\arg\max_{\pi\in\Pi}R(\pi)$.
Objective. Contextual bandit algorithms run in epochs (batches). At the start of each epoch, a new exploration (data-collection) policy is computed and is used to select arms for each round in that epoch. Consider an epoch that starts at round $\tau+1$. To construct an exploration policy $p$ for this epoch, we leverage dataset $S$ corresponding to the first $\tau$ rounds. Arguably, designing a contextual bandit algorithm boils down to designing an algorithm that uses data from prior epochs to construct the new exploration policy $p$. Hence, to simplify our analysis and focus on algorithmic insights, we study an objective that is solely a function of $p$. In particular, our objective is to construct $p$ in order to minimize the “optimal cover”, defined as $V(p,\pi^{*})$. The optimal cover is proportional to the variance of evaluating the optimal policy and has been shown to be an effective surrogate objective for pure exploration (see discussions in Krishnamurthy et al., 2023; Maurer & Pontil, 2009). [^24] At a high-level, minimizing the variance of evaluating the optimal policy allows variance-penalized policy learning algorithms to ensure better guarantees (Xu & Zeevi, 2020). Hence minimizing the optimal cover is seen as a surrogate objective for pure exploration. Furthermore, although there can be some trade-off between minimizing $V(p,\pi^{*})$ and maximizing $R(p)$, minimizing the optimal cover encourages behavior similar to the optimal policy and thus naturally results in exploration policies with high expected reward.
Input. The input dataset $S$ corresponds to data from the first $\tau$ rounds, collected under a doubling epoch schedule. [^25] That is, last $\tau/2$ rounds are collected under a single exploration policy. For any $t,t^{\prime}\in(\tau/2,\tau]$, we have $p_{t}=p_{t^{\prime}}$. For each round $t\in[\tau]$, input $S$ contains: (i) the observed context $x_{t}$, (ii) the selected arm $a_{t}$, (iii) the observed reward $r_{t}(a_{t})$, (iv) the exploration policy $p_{t}$ used at round $t$, (v) a high-probability upper bound $\alpha_{t}$ on the optimal cover ($\alpha_{t}\geq V(p_{t},\pi^{*})$), and (vi) a high-probability upper bound $M_{t}$ on the inverse probability of sampling the optimal policy under the exploration policy $p_{t}$, i.e. such that $1/p_{t}(\pi^{*}(x)|x)\leq M_{t}$. All bounds provided in dataset $S$ must simultaneously hold with probability at least $1-\delta_{\text{prior}}$, for some $\delta_{\text{prior}}\in[0,1)$. To simplify notation, we allow $t\in[\tau]$ to denote both the round index and all the corresponding data in dataset $S$. In (26), we define $\tilde{\Pi}$, a subclass of possibly optimal policies.
$\tilde{\Pi}:=\{\pi\in\Pi|p_{t}(\pi(x)|x)\geq 1/M_{t},\forall x\in\mathcal{X},t\in[\tau]\}.$ (26)
D.1 Contextual Bandit Pipeline
Algorithm 1 Contextual Bandit Pipeline: updating exploration policy at the end of an epoch
Inputs: Dataset $S$ from first $\tau$ rounds. Optimal cover guarantee for prior rounds $\{\alpha_{t}|V(p_{t},\pi^{*})\leq\alpha_{t},\forall t\in[\tau]\}$. Guarantee on sampling probability of $\pi^{*}$ in prior rounds, $\{M_{t}|p_{t}(\pi^{*}(x)|x)\geq 1/M_{t},\forall x\in\mathcal{X},t\in[\tau]\}$. Let $\tilde{\Pi}:=\{\pi\in\Pi|p_{t}(\pi(x)|x)\geq 1/M_{t},\forall x\in\mathcal{X},t\in[\tau]\}$.
Parameters: Confidence parameter $\delta\in(0,1)$. Error estimation fraction $\lambda\in(0,1/2)$. Proportional response threshold $\beta_{\max}\in(0,1)$.
1:  Split $S$: Let $S_{\text{def}}$ be data from first $\lfloor(1-\lambda)\tau\rfloor$ rounds, and $S_{\text{err}}$ be remaining data, with exploration policy $p_{\text{err}}$ and optimal cover bound $\alpha_{\text{err}}\geq V(p_{\text{err}},\pi^{*})$. We further split $S_{\text{err}}$ into $(S_{\text{err},\text{con}},S_{\text{err},\text{elim}},S_{\text{err},B})$.
2:  Estimate $(\hat{R}_{\text{elim}},\hat{\pi}_{\text{elim}})\leftarrow\text{BlackBoxPolicyEvaluator}(S_{\text{def}})$. {For any policy $\pi$, $\hat{R}_{\text{elim}}(\pi)$ is an estimate of $R(\pi)$ and can be estimated using any procedure. $\hat{\pi}_{\text{elim}}\in\tilde{\Pi}$ is ideally chosen to maximize $\hat{R}_{\text{elim}}(\cdot)$. These outputs are used for contextual arm elimination.}
3:  Estimate $\hat{U}_{\text{elim}}\leftarrow\text{CBElimErrorEstimator}(S_{\text{def}},S_{\text{err},\text{elim}},\hat{R}_{\text{elim}},\hat{\pi}_{\text{elim}},\delta,\alpha_{\text{err}})$ such that (27) holds.
$\Pr_{S_{\text{err},\text{elim}}\sim\mathcal{D}}\Big{(}\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}\Big{)}\geq 1-\delta/3.$ (27)
4:  Compute $\hat{g}\leftarrow\text{ArmEliminator}(\hat{R}_{\text{elim}},\hat{\pi}_{\text{elim}},\hat{U}_{\text{elim}},S)$ such that (28) holds.
$\hat{g}(x)\supseteq\bigg{\{}\pi(x)|\pi\in\Pi,\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi)\leq\hat{U}_{\text{elim}}\bigg{\}}\forall x.$ (28)
5:  Estimate $(\hat{f},\hat{\pi}_{\text{con}})\leftarrow\text{BlackBoxRegressor}(S_{\text{def}},S_{\text{err},\text{elim}})$. {Here $\hat{f}$ is an estimate of $f^{*}$ and can be estimated using any class/procedure. $\hat{\pi}_{\text{con}}\in\tilde{\Pi}$ is ideally chosen to maximize reward, $R_{\hat{f}}(\hat{\pi}_{\text{con}})$. These outputs are used to construct conformal arm sets.}
6:  Estimate $\hat{U}_{\text{con}},\leftarrow\text{CBConErrorEstimator}(S_{\text{def}},S_{\text{err}},\hat{f},\hat{\pi}_{\text{con}},\hat{R}_{\text{def}},\hat{\pi}_{\text{elim}},\delta,\alpha_{\text{err}})$ such that (29) holds.
$\Pr_{S_{\text{err},\text{con}}\sim\mathcal{D}}\Big{(}R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}\Big{)}\geq 1-\delta/3,$ (29)
7:  Define Conformal Arm Sets (30). Let $\text{Unif}_{\zeta}(\cdot|x)$ denote the uniform distribution over arms in $C(x,\zeta)$.
$C(x,\zeta):=\Bigg{\{}a\in\hat{g}(x):\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)\leq\frac{\hat{U}_{\text{con}}}{\zeta}\Bigg{\}}.$ (30)
8:  Select risk adjustment parameter $\eta\in[1,K]$. {Attempting to minimizes the optimal cover upper bound in Theorem D.1.}
9:  Return the next exploration policy $p_{\tau+1}$, a guarantee on its sampling probability of $\pi^{*}$, $M_{\tau+1}$, and a bound on its optimal cover $\alpha_{\tau+1}$ as defined in (31) .
$p_{\tau+1}(a|x)=(1-\beta_{\max})\text{Unif}_{\beta_{\max}/\eta}(a|x)+\int_{0}^{\beta_{\max}}\text{Unif}_{\beta/\eta}(a|x)\mbox{d}\beta.$ (31)
$M_{\tau+1}=\frac{\max_{x}|\hat{g}(x)|}{\eta\hat{U}}$
$\alpha_{\tau+1}=\mathop{\mathbb{E}}_{x\sim S_{\text{err},B}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg{]}+\frac{\max_{x}|\hat{g}(x)|}{\eta}+\sqrt{\frac{\max_{x}|\hat{g}(x)|\log(3/\delta)}{\beta_{\max}|S_{\text{err},B}|}}$
[…]
Figure 3: Visualization of data allocation in epoch $\tau$, with $\lambda=1/2$
This section describes our contextual bandit pipeline as detailed in Algorithm 1. To facilitate a truly modular contextual bandit pipeline, we employ “black box” subroutines without specific requirements on output quality whenever possible. This allows for the use of a wide range of heuristic subroutines that perform well in practice but may lack strong theoretical guarantees (e.g., methods based on deep learning or fine-tuning). To create bounds on the suboptimality of the performance of these black box subroutines, we use the error estimation theory derived above. Therefore, our algorithm begins by splitting the input dataset $S$ into $S_{\text{def}}$ and $S_{\text{err}}$,262626ensuring that the rounds in $S_{\text{err}}$ follow those in $S_{\text{def}}$. and then processes this data using several subroutines. As a result of the doubling epoch schedule, the $S_{\text{err}}$ dataset is collected under a single exploration policy $p_{\text{err}}$. We take as input a bound on the optimal cover of this exploration policy, $\alpha_{\text{err}}\geq V(p_{\text{err}},\pi^{*})$. The algorithm will be using error estimation theory in two separate instances. Additionally, we set off $S_{\text{err},B}$ to bound the optimal cover of the next exploration policy. Accordingly we split $S_{\text{err}}$ into $S_{\text{err},\text{elim}}$, $S_{\text{err},con}$, and $S_{\text{err},B}$. (see Figure 3 for outline of data allocation in round $\tau$).
We begin by applying a black box subroutine (BlackBoxPolicyEvaluator) to dataset $S_{\text{def}}$ to estimate a policy evaluator $\hat{R}_{\text{elim}}:\Pi\rightarrow[0,1]$, and construct policy $\hat{\pi}_{\text{elim}}=\operatorname*{arg\,max}_{\pi\in\tilde{\Pi}}\hat{R}_{\text{elim}}(\pi)$. Ideally, $\hat{R}_{\text{elim}}$ is equal to $R$. The closer these two policy evaluators are, the better the performance of downstream tasks.
To estimate a high-probability upper bound on the $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})$ from the output of BlackBoxPolicyEvaluator, we use a subroutine (CBElimErrorEstimator) on the dataset $S_{\text{err},\text{elim}}$. We require the output from this subroutine ($\hat{U}_{\text{elim}}$) to satisfy (27) and discuss one approach (based on error estimation) to designing such a subroutine in Section D.2 using $S_{\text{err},\text{elim}}$. With this error bound, we use subroutine (ArmEliminator) to construct $\hat{g}:\mathcal{X}\rightarrow 2^{\mathcal{A}}$ satisfying (28). Note that (28) is not a statistical requirement, and the challenge in designing such a subroutine is purely computational. A trivial way to satisfy (28) is to choose $\hat{g}(x)=\mathcal{A}$ for all $x$. We describe a more non-trivial and implementable approach to designing such a subroutine in Section D.3.
After having established a set of potentially optimal arms for every context, we apply another black box subroutine (BlackBoxRegressor) to estimate a reward model $\hat{f}:\mathcal{X}\times\mathcal{A}\rightarrow[0,1]$ and construct a policy $\hat{\pi}_{\text{con}}=\operatorname*{arg\,max}_{\pi\in\tilde{\Pi}}R_{\hat{f}}(\pi)$. Ideally, $\hat{f}$ is equal to $f^{*}$.
Once again, we describe an approach using error estimation to estimate a high probability upper bound on $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})$, a measure of the quality of $\hat{f}$ as the output of BlackBoxRegressor. We require the output ($\hat{U}_{\text{con}}$) to satisfy (29) and design a subroutine (CBConErrorEstimator) in Section D.2 using data $S_{\text{err},\text{con}}$.
Finally we can define the sets of arms we will explore in the next round. In (30) we construct conformal arm sets (CASs) for every context $x\in\mathcal{X}$. These are arms covered by potentially optimal policies according to $\hat{R}_{\text{elim}}$ (up to some error $\hat{U}_{\text{elim}}$), which are estimated to have high rewards according to $\hat{f}$ (up to some error $\hat{U}_{\text{con}}$). We then select a risk-adjustment parameter $\eta$, and construct our new exploration (data-collection) policy $p_{\tau+1}$ (see Krishnamurthy et al., 2023, for a detailed discussion on CASs and using them for designing exploration policies with low optimal cover).
Finally, we can upper bound both the optimal cover and the maximum inverse probability of sampling the arm recommended by the optimal policy (see (31)). Here we use $S_{\text{err},B}$. See Theorem D.1 for formal theoretical guarantees.
CBElimErrorEstimator (27), CBConErrorEstimator (29) and the bound for $\alpha_{\tau+1}$ (31) are the only subroutines in Algorithm 1 with a statistical guarantee.
Theorem D.1.
Consider an input dataset $S$ corresponding to data from the first $\tau$ rounds, where all bounds provided in dataset $S$ hold simultaneously with probability at least $1-\delta_{\text{prior}}$. Suppose we run Algorithm 1 on input $S$ with the following algorithmic parameters: (i) confidence parameter $\delta\in(0,1)$, (ii) error estimation fraction $\lambda\in(0,1)$, and (iii) proportional response threshold $\beta_{\max}\in(0,1)$. Let $p_{\tau+1}$, $M_{\tau+1}$, and $\alpha_{\tau+1}$ be the exploration policy returned by Algorithm 1, and bounds on its minimum sampling probability of $\pi^{*}$ and on its optimal cover respectively, as constructed in (31). Further suppose the outputs of CBElimErrorEstimator, CBConErrorEstimator and ArmEliminator satisfy the expected guarantees. That is, $\hat{U}_{\text{elim}},\hat{U}_{\text{con}}$ satisfy the high-probability guarantees in (27) and (29) respectively. Moreover, $\hat{g}$ is constructed to satisfy (28). With probability $1-\delta_{\text{prior}}-\delta$, in addition to the bounds in $S$ holding, we also have (32) holds.
$\pi^{*}(x)\in\hat{g}(x),\forall x;\;\;\;\max_{x}\frac{1}{p(\pi^{*}(x)|x)}\leq M_{\tau+1}=\frac{\max_{x}|\hat{g}(x)|}{\eta\hat{U}_{\text{con}}};$ (32)
$V(p,\pi^{*})\leq\alpha_{\tau+1}=\mathop{\mathbb{E}}_{x\sim S_{\text{err},B}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg{]}+\frac{\max_{x}|\hat{g}(x)|}{\eta}+\sqrt{\frac{\max_{x}|\hat{g}(x)|\log(3/\delta)}{\beta_{\max}|S_{\text{err},B}|}}.$
Note that with an appropriate choice of $\eta$, the optimal cover bound from Theorem D.1 primarily depends on the average size of the set $C(x,\beta_{\max}/\eta)$ – which would be small when the CBErrorEstimator outputs ($\hat{U}_{\text{elim}}$ and $\hat{U}_{\text{con}}$) are small – which in turn can only happen when outputs of the black box subroutines have fundamentally small errors.
D.2 Error Estimation for Contextual Bandits
In this section, we apply error estimation theory to develop the two CBErrorEstimator subroutines, ensuring they meet the criteria specified in equations (27) and (29). We present an overview of the terms used in Table 4. While CBElimErrorEstimatorimplements error estimation in a manner very similar to the application in Section 3.3, CBConErrorEstimatorinvolves an extra step, as the ”estimate” quantity is not directly accessible. We outline the methods and algorithms in this section, and provide detailed proofs in Section D.4.2.
Table 4: Key Error Estimation Quantities for Arm Elimination and Conformal Arm sets
Arm Elimination Contextual Arm Sets
Estimand $\theta^{\text{elim}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{elim}})$ $\theta^{\text{con}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{con}})$
$S_{\text{def}}$ Estimate $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)=\hat{R}_{\text{elim}}(\pi)-\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})$ $\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)=\mathbb{E}_{x\sim S_{err,\text{con}}}[\hat{R}_{\hat{f}}(\pi)-\hat{R}_{\hat{f}}(\hat{\pi}_{\text{con}})]$
Error $e_{\pi}^{\prime}=\theta^{\text{elim}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)$ $e_{\pi}=\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$
$S_{\text{err}}$ Estimate $\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[\theta^{\text{elim}}(\pi)]$ $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[\theta^{\text{con}}(\pi)]$
Bounding error used in Arm Elimination (27)
We begin with constructing a bound $\hat{U}_{\text{elim}}$ to satisfy (27). We define our estimand analogously to the supervised learning setting described in Section 3.3, $\theta^{\text{elim}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{elim}})$. [^27] Note that while in the setting of Section Section 3.3, $\mathcal{R}$ was a measure of risk, in this setting, $R$ represents rewards. Thus, although in Section 3.3, we had $\theta_{g}=\mathcal{R}(g_{\text{def}})-\mathcal{R}(g)$, we now have $\theta^{\text{elim}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{elim}})$, i.e. the signs have switched. This is a measure of sub-optimality of policy $\hat{\pi}_{\text{elim}}$ with respect to a particular candidate policy $\pi$. However, note that while the setting in Section 3.3 was motivated by providing a bound on the estimand $\theta_{g^{*}}$, in this setting we want to bound our estimate, $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$, based on our policy evaluator $\hat{R}_{\text{elim}}(\pi)$. We define our estimate and error as
$\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi):=\hat{R}_{\text{elim}}(\pi)-\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}}),\quad e^{\text{elim}}_{\pi}:=\theta^{\text{elim}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)$ (33)
To bound $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$ as needed in (27), we use the key insight that $\theta^{\text{elim}}(\pi^{*})=R(\pi^{*})-R(\hat{\pi}_{\text{elim}})\geq 0$ and therefore $e^{\text{elim}}_{\pi^{*}}\geq-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$. This insight is useful for two reasons. One, it allows us to use $e^{\text{elim}}_{\pi^{*}}$ as an upper bound on our quantity of interest (and hence leverage our error estimation method). Two, this allows us to satisfy 2.3, and therefore use Corollary 2.5 to improve our bounds. Now, we use $S_{\text{err},\text{elim}}$ to estimate $\theta^{\text{elim}}(\pi)$, so we define
$\hat{\theta}^{\text{elim}}_{\text{err}}(\pi):=\mathbb{E}_{(x,a,r)\sim S_{\text{err},\text{elim}}}[R(\pi)-R(\hat{\pi}_{\text{elim}})]$
to construct a quantity $\hat{u}_{\text{elim}}$ to satisfy 2.4.
$\hat{u}_{\text{elim}}(\pi,S_{\text{err},\text{elim}},\delta):=\sqrt{\frac{\log(1/\delta)}{|S_{\text{err},\text{elim}}|}}\left(\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{elim}})}+\sqrt{\alpha_{\text{err}}}\right)+\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi),$ (34)
This gives us the first bound $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})\leq e^{\text{elim}}_{\pi^{*}}\leq\hat{\xi}_{0}^{\text{elim}}(\delta):=\max_{\pi\in\tilde{\Pi}}\hat{u}_{\text{elim}}(\pi,S_{\text{err},\text{elim}},\delta)$. Now we apply localization as in Corollary 2.5 with $c=0$, set $\Pi_{0}^{\text{elim}}=\tilde{\Pi}$, and for $k\geq 1$,
$\Pi^{\text{elim}}_{k}:=\left\{\pi\in\Pi^{\text{elim}}_{k-1}|-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)\leq\hat{\xi}_{k}^{\text{elim}}(\delta)\right\},\text{ where }\hat{\xi}_{k}^{\text{elim}}(\delta):=\max_{\pi\in\Pi^{\text{elim}}_{k-1}}\hat{u}_{\text{elim}}(\pi,S_{\text{err},\text{elim}},\delta)$
By Corollary 2.5, we have (35) holds.
$\Pr_{S_{\text{err},\text{elim}}\sim D}\left(-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})\leq\hat{\xi}_{k}^{\text{elim}}(\delta)\quad\forall k\in\mathbb{N}\right)\geq 1-\delta$ (35)
Algorithm 2 CBElimErrorEstimator
Inputs: $S_{\text{def}},S_{\text{err},\text{elim}},R_{\text{elim}},\hat{\pi}_{\text{elim}},\delta$
1:  For $e^{\text{elim}}_{\pi}:=\theta^{\text{elim}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)$, $\hat{u}_{\text{elim}}$ defined in (36) satisfies 2.4.
$\hat{u}_{\text{elim}}(\pi,S_{\text{err}},\delta):=\sqrt{\frac{\log(1/\delta)}{|S_{\text{err}}|}}\left(\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{elim}})}+\sqrt{\alpha_{\text{err}}}\right)+\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi),$ (36)
2:  $\Pi_{0}^{\text{elim}}:=\{\pi\in\Pi\cup\hat{\pi}_{\text{elim}}\}$. $\hat{\xi}^{\text{elim}}_{0}(\delta):=\max_{\pi\in\Pi^{\text{elim}}_{0}}\hat{u}_{\text{elim}}(\pi,S_{\text{err}},\delta)$
3:  Now, for the localization step, we set $k=0$ and
4:  repeat
5:     $k\leftarrow k+1$
6:     $\Pi^{\text{elim}}_{k}\leftarrow\{\pi\in\Pi^{\text{elim}}_{k-1}|-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)\leq\hat{\xi}^{\text{elim}}_{k-1}(\delta)\}$
7:     $\hat{\xi}^{\text{elim}}_{k}(\delta)\leftarrow\max_{\pi\in\Pi^{\text{elim}}_{k}}\hat{u}_{\text{elim}}(\pi,S_{\text{err},\text{elim}},\delta)$
8:  until $\hat{\xi}_{k}^{\text{elim}}(\delta)\geq\hat{\xi}^{\text{elim}}_{k-1}$
9:  $\hat{U}_{\text{elim}}\leftarrow\hat{\xi}_{k}^{\text{elim}}(\delta)$.
10:  Return $\hat{U}_{\text{elim}}$.
Bounding error used in Conformal Arm Sets (29)
We now follow a similar procedure to construct a bound $\hat{U}_{\text{con}}$ to satisfy (29). In this case, our estimand is $\theta^{\text{con}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{con}})$, which again represents the sub-optimality of policy $\hat{\pi}_{\text{con}}$ with respect to a particular candidate policy $\pi$. Our goal will be to upper bound $-\theta^{\text{con}}_{\hat{f}}(\pi^{*})$, where
$\theta^{\text{con}}_{\hat{f}}(\pi)=R_{\hat{f}}(\pi)-R_{\hat{f}}(\hat{\pi}_{\text{con}})=\mathbb{E}_{x\sim D_{X}}[\hat{f}(x,\pi(x))-\hat{f}(x,\hat{\pi}_{\text{con}}(x))]$
However, $\theta^{\text{con}}_{\hat{f}}$ is not a directly estimable quantity (as $\theta^{\text{elim}}_{\hat{R}}(\pi)$ is in (33)), as it depends on the distribution over contexts $X$. Thus, we use an estimate based $S_{\text{err},\text{con}}$ for our estimate instead
$\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)=\mathbb{E}_{x\sim S_{\text{err},\text{con}}}[\hat{f}(x,\pi(x))-\hat{f}(x,\hat{\pi}_{\text{con}}(x))],\quad e^{\text{con}}_{\pi}:=\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$ (37)
Again, as $\theta^{\text{con}}(\pi^{*})=R(\pi^{*})-R(\hat{\pi}_{\text{con}})\geq 0$, we know that $e_{\text{con},\pi^{*}}\geq-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi^{*})$. As in (34), we construct $\hat{u}_{\text{con}}(\pi,S_{\text{err},\text{con}},\delta)$.
$\hat{u}_{\text{con}}(\pi,S_{\text{err},\text{con}},\delta)$ $:=\sqrt{\frac{\log(1/\delta)}{|S_{\text{err},\text{con}}|}}\left(\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{con}})}+\sqrt{\alpha_{\text{err}}}\right)+\hat{\theta}^{\text{con}}_{\text{err}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi),$ (38)
We then localize over the classes
$\Pi^{\text{con}}_{k}:=\left\{\pi\in\Pi^{\text{con}}_{k-1}|-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)\leq\hat{\xi}_{k}^{\text{con}}(\delta)\right\},\text{ where }\hat{\xi}_{k}^{\text{con}}(\delta):=\max_{\pi\in\Pi^{\text{con}}_{k-1}}\hat{u}_{\text{con}}(\pi,S_{\text{err},\text{con}},\delta)$
By Corollary 2.5, we have (39) holds, this time we set the probability to be $21-\delta/$ 2.
$\Pr_{S_{\text{err},\text{con}}\sim D}\left(-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi^{*})\leq\hat{\xi}_{k}^{\text{con}}(\delta/2)\quad\forall k\in\mathbb{N}\right)\geq 1-\delta/2$ (39)
Finally, to bound our true quantity of interest, we use Hoeffding’s inequality again,
$\Pr_{S_{\text{err},\text{con}}\sim D}\left(-\theta^{\text{con}}_{\hat{f}}(\pi^{*})\leq-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi^{*})+\sqrt{\frac{2\log(2/\delta)}{|S_{\text{err},\text{con}}|}}\right)\geq 1-\delta/2$ (40)
See details in Section D.4.2.
Algorithm 3 CBConErrorEstimator
Inputs: $S_{\text{def}},S_{\text{err}},\hat{f},\hat{\pi}_{\text{con}},\delta$
1:  For $e^{\text{con}}_{\pi}:=\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$, $\hat{u}_{\text{con}}$ defined in (41) satisfies 2.4.
$\hat{u}_{\text{con}}(\pi,S_{\text{err}},\delta):=\sqrt{\frac{\log(1/\delta)}{|S_{\text{err}}|}}\left(\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{con}})}+\sqrt{V(p_{\text{err}},\pi)}\right)+\hat{\theta}^{\text{con}}_{\text{err}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi),$ (41)
2:  $\Pi^{\text{con}}_{0}:=\{\pi\in\Pi\cup\hat{\pi}_{\text{con}}\}$. $\hat{\xi}^{\text{con}}_{0}(\delta):=\max_{\pi\in\Pi^{\text{con}}_{0}}\hat{u}_{\text{con}}(\pi,S_{\text{err},\text{con}},\delta/3)$
3:  Now, for the localization step, we set $k=0$ and
4:  repeat
5:     $k\leftarrow k+1$
6:     $\Pi^{\text{con}}_{k}\leftarrow\{\pi\in\Pi^{\text{con}}_{k-1}|-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)\leq\hat{\xi}^{\text{con}}_{k-1}(\delta)\}$
7:     $\hat{\xi}^{\text{con}}_{k}(\delta)\leftarrow\max_{\pi\in\Pi_{k}}\hat{u}_{\text{con}}(\pi,S_{\text{err},\text{con}},\delta)$
8:  until $\hat{\xi}^{\text{con}}_{k}(\delta)\geq\hat{\xi}^{\text{con}}_{k-1}(\delta)$
9:  $\hat{U}_{\text{con}}\leftarrow\hat{\xi}^{\text{con}}_{k}(\delta/2)$.
10:  Return $\hat{U}_{\text{con}}+\sqrt{\frac{2\log(2/\delta)}{|S_{\text{err},\text{con}}|}}$.
Corollary D.2.
Given $S_{\text{err}},S_{\text{def}},R_{\text{elim}},\hat{\pi}_{\text{elim}},\delta$, (43) holds.
$\Pr_{S_{\text{err},\text{elim}}\sim\mathcal{D}}\left(\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{\xi}^{\text{elim}}_{k}(\delta)\quad\forall k\in\mathbb{N}\right)\geq 1-\delta,$ (42)
Independently, given $S_{\text{err}},S_{\text{def}},\hat{f},\hat{\pi}_{\text{con}},\delta$, (42) holds.
$\Pr_{S_{\text{err},\text{con}}\sim\mathcal{D}}\left(R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{\xi}^{\text{con}}_{k}(\delta/2)+\sqrt{\frac{2\log(2/\delta)}{|S_{\text{err},\text{con}}|}}\quad\forall k\in\mathbb{N}\right)\geq 1-\delta,$ (43)
D.3 Contextual Arm Elimination
In this section we construct a simple Arm Eliminator which uses an upper confidence bound (UCB) style argument to guess which arms to include, then checks that no arms which perform well enough under $\hat{R}_{\text{elim}}$ have been excluded. If there exist such arms, ArmEliminator includes more arms.
Algorithm 4 ArmEliminator
Inputs: $\hat{R}_{\text{elim}},\hat{\pi}_{\text{elim}},\hat{U}_{\text{elim}},S$.
1:  Estimate $\hat{f}_{\text{mean}},\hat{f}_{0pt}\leftarrow\text{BlackBoxCIEstimator}(S)$
[…]
3:  while not ValidEliminator do
4:     Set $\hat{g}(x)\leftarrow\{a|\hat{f}_{\text{mean}}(x,a)+\gamma*\hat{f}_{0pt}(x,a)\geq\max_{a}(\hat{f}_{\text{mean}}(x,a)-\gamma*\hat{f}_{0pt}(x,a))\}$ for all $x$.
[…]
6:     $\gamma\leftarrow 2*\gamma$.
7:  end while
8:  Return $\hat{g}$.
Where ValidateEliminator is designed to output (44), which we approximate with (45).
$\mathbf{1}\Bigg{(}\min_{\{\pi\in\tilde{\Pi}|\pi(x)\notin g(x)\forall x\}}\hat{R}_{\text{elim}}(\pi)-\hat{R}_{\text{elim}}(\pi^{*})>\hat{U}_{\text{elim}}\Bigg{)}$ (44)
$\approx\mathbf{1}\Bigg{(}\min_{\{\pi\in\tilde{\Pi}|\pi(x)\notin g(x)\forall x\in S\}}\hat{R}_{\text{elim}}(\pi)-\hat{R}_{\text{elim}}(\pi^{*})>\hat{U}_{\text{elim}}\Bigg{)}=\text{ValidateEliminator}(\hat{R}_{\text{def}},\hat{\pi}_{\text{def}},\hat{U}^{\prime},S).$ (45)
D.4 Details for Contextual Bandit Pipeline
In this Appendix we prove Theorem D.1 for the Contextual Bandit pipeline and formalize the error estimation argument in the Contextual Bandit setting.
D.4.1 Proof of Theorem 2
We begin relating conformal arm sets to classical notions in conformal prediction ((Vovk et al., 2005)).
Lemma D.3 (Conformal Uncertainty).
Suppose $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$, we then have:
$\Pr_{x\sim D_{\mathcal{X}}}(\pi^{*}(x)\in C(x,\zeta))\geq 1-\zeta,\;\;\;\forall\zeta\in(0,1].$ (46)
Proof.
Suppose $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$, we then have (47) holds.
$\Pr_{x\sim D_{\mathcal{X}}}(\pi^{*}(x)\notin C(x,\zeta))$ (47)
=$\stackrel{{\scriptstyle(i)}}{{=}}\Pr_{x\sim D_{\mathcal{X}}}\Big{(}\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,\pi^{*}(x))>\frac{\hat{U}_{\text{con}}}{\zeta}\Big{)}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}\frac{\mathop{\mathbb{E}}_{x\sim D_{\mathcal{X}}}[\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,\pi^{*}(x))]}{\hat{U}_{\text{con}}/\zeta}\stackrel{{\scriptstyle(iii)}}{{=}}\frac{R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})}{\hat{U}_{\text{con}}/\zeta}\stackrel{{\scriptstyle(iv)}}{{\leq}}\zeta.$
Here (i) follows from (30), (ii) follows from Markov’s inequality, (iii) follows by definition of $R_{\hat{f}}$ (25), and (iv) follows from $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$. ∎
Now, we formalize the intuition that if an arm $a$ is likely to be a good arm for context $x$ (either by belonging to both $\hat{g}(x)$ and $C(x,\beta_{\max}/\eta)$, or just to $\hat{g}(x)$), then the learner has a positive probability of pulling arm $a$ under exploration policy $p$.
Lemma D.4.
We have (48) holds.
$p(a|x)$$\geq\begin{cases}\frac{1-\beta_{\max}}{|C(x,\beta_{\max}/\eta)|}+\frac{\beta_{\max}}{|\hat{g}(x)|},&\text{if $a\in C(x,\beta_{\max}/\eta)$}\\ \frac{\eta}{|\hat{g}(x)|}\frac{\hat{U}_{\text{con}}}{\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)},&\text{if $a\notin C(x,\beta_{\max}/\eta)$ and $a\in\hat{g}(x)$}\\ 0,&\text{otherwise.}\end{cases}$ (48)
$\geq\begin{cases}\frac{1}{|\hat{g}(x)|},&\text{if $a\in C(x,\beta_{\max}/\eta)$}\\ \frac{\eta\hat{U}_{\text{con}}}{|\hat{g}(x)|},&\text{if $a\notin C(x,\beta_{\max}/\eta)$ and $a\in\hat{g}(x)$}\\ 0,&\text{otherwise.}\end{cases}$
Proof.
Recall that $\text{Unif}_{\zeta}(\cdot|x)$ denotes the uniform distribution over arms in $C(x,\zeta)$ and $p$ is given by (49).
$p(a|x)=(1-\beta_{\max})\text{Unif}_{\beta_{\max}/\eta}(a|x)+\int_{0}^{\beta_{\max}}\text{Unif}_{\beta/\eta}(a|x)\mbox{d}\beta.$ (49)
Clearly if $a\notin\hat{g}(x)$, we have $p(a|x)=0$. The rest of our analysis to lower bound $p(a|x)$ is divided into two cases based on whether $a$ lies in $C(x,\beta_{\max}/\eta)$.
Case 1 ($a\in C(x,\beta_{\max}/\eta)$). Note that $C(x,\beta_{\max}/\eta)\subseteq C(x,\beta/\eta)\subseteq\mathcal{A}$ for all $\beta\in[0,\beta_{\max}]$. Hence, $a\in C(x,\beta/\eta)$ for all $\beta\in[0,\beta_{\max}]$. Therefore, in this case, $p(a|x)\geq\frac{1-\beta_{\max}}{|C(x,\beta_{\max}/\eta)|}+\frac{\beta_{\max}}{|\hat{g}(x)|}\geq\frac{1}{|\hat{g}(x)|}$.
Case 2 ($a\notin C(x,\beta_{\max}/\eta)$ and $a\in\hat{g}(x)$). For this case, the proof follows from (50).
$p(a|x)\geq$ $\int_{0}^{\beta_{\max}}\frac{I[a\in C(x,\beta/\eta)]}{|C(x,\beta/\eta)|}\mbox{d}\beta$ (50)
$\stackrel{{\scriptstyle(i)}}{{\geq}}$ $\frac{1}{|\hat{g}(x)|}\int_{0}^{\beta_{\max}}I[a\in C(x,\beta/\eta)]\mbox{d}\beta=\frac{1}{|\hat{g}(x)|}\int_{0}^{\beta_{\max}}I[a\in C(x,\beta/\eta)]\mbox{d}\beta$
$\stackrel{{\scriptstyle(ii)}}{{\geq}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\int_{0}^{\beta_{\max}}I[a\in C(x,\beta/\eta)]\mbox{d}\beta$
=$\stackrel{{\scriptstyle(iii)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\int_{0}^{1}I[a\in C(x,\beta/\eta)]\mbox{d}\beta$
=$\stackrel{{\scriptstyle(iv)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\int_{0}^{1}I\bigg{[}\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)\leq\frac{\eta\hat{U}_{\text{con}}}{\beta}\bigg{]}\mbox{d}\beta$
$=$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\int_{0}^{1}I\bigg{[}\beta\leq\frac{\eta\hat{U}_{\text{con}}}{\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)}\bigg{]}\mbox{d}\beta$
=$\stackrel{{\scriptstyle(v)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\frac{\eta\hat{U}_{\text{con}}}{\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)}$
$\stackrel{{\scriptstyle(vi)}}{{\geq}}$ $\frac{\eta\hat{U}_{\text{con}}I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}$
where (i) is because the size of a set’s intersection with $\hat{g}(x)$ can be no larger than the size of $\hat{g}(x)$; (ii) follows from $I(a\notin C(x,\beta_{\max}/\eta))\leq 1$; (iii) follows from the fact that if $a\notin C(x,\beta_{\max}/\eta)$ then $a\notin C_{m}(x,\beta/\eta)$ for all $\beta\geq\beta_{\max}$; (iv) follows from (30); (v) follows from the fact that $\frac{\eta\hat{U}_{\text{con}}}{\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)}\leq 1$ when $a\notin C(x,\beta_{\max}/\eta)$; and (vi) follows from the fact that $\hat{f}:\mathcal{X}\times\mathcal{A}\rightarrow[0,1]$. ∎
Now, in Lemma D.5, we use the bounds derived in Lemma D.4 to establish bounds on the optimal cover $V(p,\pi^{*})$.
Lemma D.5.
Suppose $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$ and $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$. Further suppose (28) holds, therefore $\pi^{*}(x)\in\hat{g}(x)$ for all $x$. We then have (51) holds.
$V(p,\pi^{*})$ $\leq\mathop{\mathbb{E}}_{x\sim D_{\mathcal{X}}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}+\frac{|\hat{g}(x)|}{\eta}\frac{(\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,\pi^{*}(x)))}{\hat{U}_{\text{con}}}\Bigg{]}$ (51)
$\leq\mathop{\mathbb{E}}_{x\sim D_{\mathcal{X}}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg{]}+\frac{\max_{x}|\hat{g}(x)|}{\eta}.$
Proof.
From Lemma D.4, we have (52) holds.
$\frac{I(\pi^{*}(x)\in C(x,\beta_{\max}/\eta))}{p(\pi^{*}(x)|x)}\leq\frac{1}{\frac{1-\beta_{\max}}{|C(x,\beta_{\max}/\eta)|}+\frac{\beta_{\max}}{|\hat{g}(x)|}}\leq\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}},$ (52)
and $\frac{I(\pi^{*}(x)\notin C(x,\beta_{\max}/\eta))}{p(\pi^{*}(x)|x)}\leq\frac{|\hat{g}(x)|}{\eta}\frac{\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,\pi^{*}(x))}{\hat{U}_{\text{con}}}.$
We now bound the optimal cover as follows,
$V(p,\pi^{*})=\mathop{\mathbb{E}}_{x\sim D_{\mathcal{X}}}\bigg{[}\frac{1}{p(\pi^{*}(x)|x)}\bigg{]}$ (53)
$=$ $\mathop{\mathbb{E}}_{x\sim D_{\mathcal{X}}}\bigg{[}\frac{I[\pi^{*}(x)\in C(x,\beta_{\max}/\eta)]+I[\pi^{*}(x)\notin C(x,\beta_{\max}/\eta)]}{p(\pi^{*}(x)|x)}\bigg{]}$
$\stackrel{{\scriptstyle(i)}}{{\leq}}$ $\mathop{\mathbb{E}}_{x\sim D_{\mathcal{X}}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}+\frac{|\hat{g}(x)|}{\eta}\frac{(\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,\pi^{*}(x)))}{\hat{U}_{\text{con}}}\Bigg{]}$
$\stackrel{{\scriptstyle(ii)}}{{\leq}}$ $\mathop{\mathbb{E}}_{x\sim D_{\mathcal{X}}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg{]}+\frac{\max_{x}|\hat{g}(x)|}{\eta}.$
Here (i) follows from from (52) and (ii) follows from $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$. ∎
Finally, we combine results from the Lemmas above to prove Theorem D.1. See D.1
Proof.
Since the bounds in $S$ hold with probability at least $1-\delta_{\text{prior}}$ and under this event (27) holds, we have: the bounds in $S$, $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$, and $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$ hold with probability at least $31-\delta_{\text{prior}}-2\delta/$ 3. Hence, by the union bound, it is sufficient to prove (32) with probability at least $31-\delta/$ 3 under these events. Therefore moving forward, we assume $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$ and $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$ hold.
Now note that $\hat{g}$ is computed in a way that satisfies (28). Since $\hat{R}_{\text{elim}}(\hat{\pi}_{\\ elim})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$, we therefore have $\hat{g}(x)\in\pi^{*}(x)$ for all $x$. Hence from $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$, $\hat{g}(x)\in\pi^{*}(x)$ for all $x$, and Lemmas D.3, D.4 and D.5, we have (54) holds.
$\pi^{*}(x)\in\hat{g}(x),\forall x;\;\;\;\max_{x}\frac{1}{p(\pi^{*}(x)|x)}\leq\frac{\max_{x}|\hat{g}(x)|}{\eta\hat{U}_{\text{con}}};$ (54)
$V(p,\pi^{*})\leq\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{x}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg{]}+\frac{\max_{x}|\hat{g}(x)|}{\eta}$
Finally, by Hoeffding’s inequality, with probability at least $31-\delta/$ 3, we have
$V(p,\pi^{*})$ $\leq\mathop{\mathbb{E}}_{x\sim\mathcal{D}_{x}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg{]}+\frac{\max_{x}|\hat{g}(x)|}{\eta}$
$\leq\mathop{\mathbb{E}}_{x\sim S_{\text{err},B}}\Bigg{[}\frac{|C(x,\beta_{\max}/\eta)|}{1-\beta_{\max}+\beta_{\max}\frac{|C(x,\beta_{\max}/\eta)|}{|\hat{g}(x)|}}\Bigg{]}+\sqrt{\frac{\max_{x}|\hat{g}(x)|\log(3/\delta)}{\beta_{\max}|S_{\text{err},B}|}}+\frac{\max_{x}|\hat{g}(x)|}{\eta}.$
D.4.2 Proof of Error Estimation for Contextual Bandits
In this section, we present the steps required to apply the Error Estimation method to the Contextual Bandit setting, and ultimately prove Corollary D.2.
Proof.
Recall that we have two subroutines:
  • CBElimErrorEstimator constructs $\hat{U}_{\text{elim}}$ as a bound on the error in the Policy Evaluation step, such that (27) holds:
    $\Pr_{S_{\text{err},\text{elim}}\sim\mathcal{D}}\left(\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}\right)\geq 1-\delta/3$
  • CBConErrorEstimator constructs $\hat{U}_{\text{con}}$ as a bound on the error in the Reward model step, such that (29) holds:
    $\Pr_{S_{\text{err},\text{con}}\sim\mathcal{D}}\Big{(}R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}\Big{)}\geq 1-\delta/3$
Arm Elimination Contextual Arm Sets
Estimand $\theta^{\text{elim}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{elim}})$ $\theta^{\text{con}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{con}})$
$S_{\text{def}}$ Estimate $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)=\hat{R}_{\text{elim}}(\pi)-\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})$ $\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)=\mathbb{E}_{x\sim S_{err,\text{con}}}[\hat{R}_{\hat{f}}(\pi)-\hat{R}_{\hat{f}}(\hat{\pi}_{\text{con}})]$
Error $e_{\pi}^{\prime}=\theta^{\text{elim}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)$ $e_{\pi}=\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$
$S_{\text{err}}$ Estimate $\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[\theta^{\text{elim}}(\pi)]$ $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[\theta^{\text{con}}(\pi)]$
Recall the Key Error Estimation Quantities for Arm Elimination and Conformal Arm sets in Table 4.
Step 1: Bounding the error in Policy Evaluation. Recall that we want to bound $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$, and since $\theta^{\text{elim}}(\pi^{*})=R(\pi^{*})-R(\hat{\pi}_{\text{elim}})\geq 0$ (recall we define $\pi^{*}=\arg\max_{\pi\in\Pi}R(\pi)$, and construct $\Pi$ to include $\hat{\pi}_{\text{elim}}$), it suffices to find an upper bound for $e^{\text{elim}}_{\pi^{*}}$. Using the error estimation heuristic, we introduce $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)$ as in Table 4, and write
$e^{\text{elim}}_{\pi}$ $=\theta^{\text{elim}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)+\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)-\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)$
$=\underbrace{\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)}_{A}+\underbrace{\theta^{\text{elim}}(\pi)-\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)}_{B}$
  1. 1.
    We can find $A$ directly as both terms are estimable. Note that we estimate the first term using inverse propensity scores. $A=\mathbb{E}_{(a,x)\sim S_{\text{err}}}\left[\frac{\mathbf{1}_{a=a^{\prime}}r(a^{\prime},x)(\pi(a^{\prime}|x)-\hat{\pi}_{\text{elim}}(a^{\prime}|x)}{p_{\text{err}}(a^{\prime}|x)}\right]-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)$.
  2. 2.
    For $B$, note that this term corresponds to $e^{\text{elim}}_{\text{err},\pi}$ from the general framework introduce in Section 3. We now construct a bound that satisfies 2.4. In particular, by Freedman’s inequality (Dudik et al., 2011), we have for any $\pi\in\tilde{\Pi}$
    $B\leq\sqrt{\frac{\log(1/\delta)}{n}}\left(\sqrt{V(p_{\text{err}},\pi)}+\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{elim}})}\right)$ (55)
    with probability at least $1-\delta$
Combining these terms, we have
$e^{\text{elim}}_{\pi}\leq\sqrt{\frac{\log(1/\delta)}{n}}\left(\sqrt{V(p_{\text{err}},\pi)}+\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{elim}})}\right)+\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)$
Now we define
$\hat{u}_{\text{elim}}(\pi,S_{\text{err}},\delta):=\sqrt{\frac{\log(1/\delta)}{n}}\left(\sqrt{\alpha_{\text{err}}}+\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{elim}})}\right)+\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)$ (56)
Note that $\hat{u}_{\text{elim}}(\pi,S_{\text{err}},\delta)$ satisfies 2.4 rather than 2.1, as we use $\alpha_{\text{err}}\geq V(p_{\text{err}},\pi^{*})$ for the specific instance $\pi^{*}$. Thus, we have the initial bound
$e_{\pi^{*}}^{\text{elim}}\leq\hat{u}_{\text{elim}}(\pi^{*},S_{\text{err}},\delta)\leq\max_{\pi\in\Pi}\left(\hat{u}_{\text{elim}}(\pi,S_{\text{err}},\delta)\right)$ (57)
Step 2: Improving the error bound in Policy Evaluation using localization. Now, given $\theta^{\text{con}}(\pi^{*})\geq 0$, we have satisfied 2.3, and can thus apply Corollary 2.5. Let $\Pi^{\text{elim}}_{0}=\tilde{\Pi}$, and define
$\hat{\xi}^{\text{elim}}_{k}(\delta):=\max_{\pi\in\Pi^{\text{elim}}_{k-1}}\hat{u}_{\text{elim}}(\pi,S_{\text{err}},\delta),\;\;\;\Pi^{\text{elim}}_{k}:=\left\{\pi\in\Pi^{\text{elim}}_{k-1}|-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)\leq\xi_{k}\right\},$
Recalling that $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})\leq e^{\text{elim}}_{\pi^{*}}$, this gives us the final bound
$\Pr_{S_{\text{err},\text{elim}}\sim\mathcal{D}}\left(-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})\leq\hat{\xi}^{\text{elim}}_{k}(\delta)\forall k\right)\geq 1-\delta.$
Step 3: Bounding the error in the reward model. For the reward model, recall that we are interested in bounding $-\theta^{\text{con}}_{\hat{f}}(\pi^{*})$, where
$\theta^{\text{con}}_{\hat{f}}(\pi)=R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi)=\mathbb{E}_{x\sim\mathcal{D}}[\hat{f}(x,\hat{\pi}_{\text{con}}(x)-\hat{f}(x,\pi(x))].$
However, unlike the quantity in arm elimination, this is not directly computable as it an average over an unknown distribution of contexts. Thus we will instead use the estimate
$\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[R(\pi)-R(\hat{\pi}_{\text{con}})],$
and use error estimation to upper bound the maximum error, and then use a single error estimate to bound the difference between this estimate and our true estimate of interest. This case is now almost identical to the arm elimination case. We similarly define $e^{\text{con}}_{\pi}=\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$. Since $\theta^{\text{con}}(\pi^{*})(=R(\pi)-R(\hat{\pi}_{\text{con}}))\geq 0$, it suffices to find an upper bound for $e^{\text{con}}_{\pi^{*}}$ to bound $-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi^{*})$.
$e_{\pi}$ $=\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)+\hat{\theta}^{\text{con}}_{\text{err}}(\pi)-\hat{\theta}^{\text{con}}_{\text{err}}(\pi)$
$=\underbrace{\hat{\theta}^{\text{con}}_{\text{err}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)}_{A}+\underbrace{\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\text{err}}(\pi)}_{B}$
  1. 1.
    Estimating $\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$ directly and using an unbiased estimate to evaluate $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)$.
    $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)$ $=\hat{R}^{\text{err}}(\pi)-\hat{R}^{\text{err}}(\hat{\pi}_{\text{con}})$
    $=\mathbb{E}_{(a,x)\sim S_{\text{err}}}\left[\sum_{a^{\prime}\in\mathcal{A}}\frac{\mathbf{1}_{a=a^{\prime}}r(a^{\prime},x)(\pi(a^{\prime}|x)-\hat{\pi}_{\text{con}}(a^{\prime}|x)}{p_{\text{err}}(a^{\prime}|x)}\right]$
    So
    $A=\mathbb{E}_{(a,x)\sim S_{\text{err}}}\left[\sum_{a^{\prime}\in\mathcal{A}}\frac{\mathbf{1}_{a=a^{\prime}}r(a^{\prime},x)(\pi(a^{\prime}|x)-\hat{\pi}_{\text{con}}(a^{\prime}|x)}{p_{\text{err}}(a^{\prime}|x)}\right]-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$
  2. 2.
    Note that we can write
    $\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\text{err}}(\pi)$ $=R(\pi)-\hat{R}^{\text{err}}(\pi)+\hat{R}^{\text{err}}(\hat{\pi}_{\text{\text{con}}})-R(\hat{\pi}_{\text{\text{con}}})$
    From Freedman’s inequality (Dudik et al., 2011), we have with probability at least $21-\delta/$ 2
    $B\leq\sqrt{\frac{\log(2/\delta)}{n}}\left(\sqrt{V(p_{\text{err}},\pi)}+\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{con}})}\right)$
Combining the terms, we have that for any given policy $\pi$, with probability at least $1-\delta$.
$e^{\text{con}}_{\pi}\leq\sqrt{\frac{\log(1/\delta)}{n}}\left(\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{con}})}+\sqrt{V(p_{\text{err}},\pi)}\right)+\hat{\theta}^{\text{con}}_{\text{err}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$
Now, as above, we define $\hat{u}_{\text{con}}(\pi,S_{\text{err}},\delta)$ to satisfy 2.4.
$\hat{u}_{\text{con}}(\pi,S_{\text{err}},\delta):=\sqrt{\frac{\log(1/\delta)}{n}}\left(\sqrt{V(p_{\text{err}},\hat{\pi}_{\text{con}})}+\sqrt{\alpha_{\text{err}}}\right)+\hat{\theta}^{\text{con}}_{\text{err}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$ (58)
Now, since $\theta^{\text{con}}(\pi^{*})\geq 0$, we have satisfied 2.3 and can apply Corollary 2.5. We note let $\Pi^{\text{con}}_{0}=\tilde{\Pi}$, and define
$\hat{\xi}^{\text{con}}_{k}(\delta/2):=\max_{\pi\in\Pi^{\text{con}}_{k-1}}\hat{u}_{\text{con}}(\pi,S_{\text{err},\text{con}},\delta/2),\;\;\;\Pi^{\text{con}}_{k}:=\left\{\pi\in\Pi^{\text{con}}_{k-1}|-\theta^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)\leq\hat{\xi}_{k}(\delta/2)\right\},$
This gives us
$\Pr_{S_{\text{err},\text{con}}\sim\mathcal{D}}\left(-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi^{*})\leq\hat{\xi}^{\text{con}}_{k}(\delta/2)\forall k\right)\geq 1-\delta/2.$
Finally, to account for having used the estimate $\hat{\theta}_{\hat{f},\text{err}_{X}}(\pi)$ rather than $\theta_{\hat{f}}(\pi)$, we bound the difference by recognizing that $\hat{f}$ is determined independently of $S_{\text{err}}$, and we can therefore use a classic Hoeffding inequality, along with a union bound argument to find that with probability at least $21-\delta/$ 2,
$\Pr_{S_{\text{err},\text{con}}\sim\mathcal{D}}\left(-\theta^{\text{con}}_{\hat{f}}(\pi^{*})\leq\hat{\xi}^{\text{con}}_{k}(\delta/2)\forall k+\sqrt{\frac{2\log(2/\delta)}{|S_{\text{err},\text{con}}|}}\right)\geq 1-\delta.$